### v1

# Robustness of topological phases on aperiodic lattices

Abstract.
We study the robustness of topological phases on aperiodic lattices by constructing *-homomorphisms from the groupoid model to the coarse-geometric model of observable C*-algebras. These *-homomorphisms induce maps in K-theory and Kasparov theory. We show that the strong topological phases in the groupoid model are detected by position spectral triples. We show that topological phases coming from stacking along another Delone set are always weak in the coarse-geometric sense.
1. Introduction
In this article, we aim at a new look at the following question:
How to understand the robustness of topological phases of a discrete physical system, described by a generic aperiodic point pattern?
We begin with some background around it. Topological insulators are materials that are insulating in the bulk but nevertheless possess metallic edges. The current flowing on the edge is usually quite robust under disorder coming from impurity of the crystal. Starting from Bellissard, van Elst and Schulz-Baldes [Bellissard-van_Elst-Schulz_Baldes:NCG_of_QHE, Bellissard:K-theory_in_solid_state_physics], the non-commutative and C*-algebraic tools have been intensively applied to the study of such materials and in particular, leading to an interpretation of the Kubo formula in the integer quantum Hall effect (IQHE) as a quantised pairing between K-theory and cyclic cohomology.
In physics literature, the model Hamiltonian of a topological insulator is usually given by a tight-binding, short-range operator supported on a periodic, square lattice, invariant under the translation by its unit cell vector. Such a Hamiltonian $H$ (or its spectral projection) belong to a (noncommutative) torus $A_{\vartheta}=\mathbb{C}\rtimes_{\vartheta}\mathbb{Z}^{d}$, the noncommutativity coming from a 2-cocycle twist $\vartheta$ given by the external magnetic field. If the Hamiltonian is not translation invariant, then Bellissard suggested to replace $A_{\vartheta}$ by a twisted crossed product $\mathrm{C}(\Omega)\rtimes_{\vartheta}\mathbb{Z}^{d}$, where $\Omega$ is a compact space called the hull of $H$. This description is applicable even if the underlying space is no longer a square lattice, but comes from a tiling subject to some properties, cf. [Anderson-Putnam:Substitution_tilings, Sadun-Williams:Tilings_Cantor_fiber_bundles].
There are, however, amorphous materials, like liquid crystals and glass, that cannot be modelled using such methods. Even if for quasi-crystals such that the method of [Anderson-Putnam:Substitution_tilings, Sadun-Williams:Tilings_Cantor_fiber_bundles] applies, then one has to essentially provide a (non-canonical) $\mathbb{Z}^{d}$-labelling of the underlying lattice, under which the Hamiltonian remains short-range (cf.[Bourne-Prodan:Noncommutative_Chern_numbers]).
These suggest to use more general point sets in $\mathbb{R}^{d}$, modelling the atomic sites of this material, as their underlying geometric spaces. The dynamics thereon are studied by more general actions other than a $\mathbb{Z}^{d}$-translation. It was explained in [Bellissard-Herrmann-Zarrouati:Hulls_aperiodic_solids] that such point sets should be Delone sets (Definition 2.14). Let $\Lambda$ be a Delone set, whose points are considered as the sites in a physical system. How should one model the observable C*-algebra $A$ from it?
The choice of an observable C*-algebra should be aligned with the following principle: it should be large enough to contain all possible Hamiltonians, and small enough to supply useful homotopy theory (K-theory). In recent years, there have been two main approaches to this modelling problem, which provide toolkits to compute numerical invariants of topological phases:
  • “dynamical” approach describes a crossed product C*-algebra, covariant for the groupoid actions on the aperiodic point pattern, then restricts to the dynamical hull of the point pattern. This gives a groupoid C*-algebra.
  • “universal” approach describes a C*-algebra which is stable under all short-range, locally–finite-rank perturbations. This leads to a Roe C*-algebra.
The groupoid approach gives an étale groupoid $\mathcal{G}_{\Lambda}\rightrightarrows\Omega_{0}$ from a Delone set $\Lambda$, and yields a “tight-binding” groupoid C*-algebra $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$ , cf. [Bellissard-Herrmann-Zarrouati:Hulls_aperiodic_solids, Bourne-Prodan:Noncommutative_Chern_numbers, Bourne-Mesland:Aperiodic, Mesland-Prodan:Interacting]. This is a generalisation of the well-studied periodic model: if $\Lambda=\mathbb{Z}^{d}$, then $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\simeq\mathrm{C}^{*}(\mathbb{Z}^{d})$ is the group C*-algebra of $\mathbb{Z}^{d}$. Here tight-binding means the following: the regular representation of $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$ is given by a Hilbert $\mathrm{C}(\Omega_{0})$-module, or equivalently, a continuous field of Hilbert spaces over $\Omega_{0}$, where $\Omega_{0}$ is unit space of the groupoid $\mathcal{G}_{\Lambda}$. The fibres of this continuous field are canonically identified with $\ell^{2}(\omega)$, where $\omega\in\Omega_{0}$ is a Delone set, either as a translated copy of $\Lambda$, or as a weak*-limit of such sets. Thus $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$ consists of families of model Hamiltonians $(H_{\omega})_{\omega\in\Omega_{0}}$, where $H_{\omega}$ acts on $\ell^{2}(\omega)$, in a covariant way with respect to the groupoid action.
The coarse-geometric approach describes a C*-algebra which is stable under all possible short-range perturbations. This gives rise to a (uniform or non-uniform) Roe C*-algebra $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$ from $\Lambda$, cf. [Kubota:Controlled_topological_phases, Ewert-Meyer:Coarse_geometry]. We choose the non-uniform Roe C*-algebras, whose advantages over the uniform ones were explained in [Ewert-Meyer:Coarse_geometry].
In this setup, a Delone set $\Lambda\subseteq\mathbb{R}^{d}$ is considered as a discrete metric space. Then $L^{2}(\mathbb{R}^{d})$ carries an ample representation of $\mathrm{C}_{0}(\Lambda)$, which generates a Roe C*-algebra $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$. Physically, a short-range Hamiltonian $H$ has matrix coeffcients $H_{x,y}\mathrel{\vcentcolon=}\left\langle x\,\middle|H\middle|\,\mathopen{}y\right\rangle$ of fast enough decay, thus can be approximated by controlled operators. Every unit cell should have finite degrees of freedom, coming from the number of electron orbits and their spins. Thus the restriction of $H$ to any finite region should have finite-rank. Therefore, the Roe C*-algebra $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$ can be viewed as the universal C*-algebra that contains all such Hamiltonians.
Topological phases on materials are represented by a K-theory class of the C*-algebra of observables, constructed from the Hamiltonian $H$. If the system consists of conjugate-linear symmetries, then the C*-algebra is replaced by a $\mathbb{Z}/2$-graded “real” C*-algebra of the form $A_{\mathbb{C}}\otimes\mathbb{C}\ell_{p,q}$, or its corresponding real subalgebra $A_{\mathbb{R}}\otimes\mathrm{C}\ell_{p,q}$. We wish to describe the topological phases with their numerical invariants. That is, we seek for maps
(1.1) $\operatorname{KK}(\mathbb{R},A\otimes\mathrm{C}\ell_{p,q})\to\mathbb{R}$
generating a real number from a prescribed topological phase. Moreover, we wish to understand the robustness of such numbers.
If we model a topological insulator on a Delone set $\Lambda$ by the Roe C*-algebra $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$, then its topological phase is robust in a very strong sense: such topological phases (and hence their numerical invariants as in (1.1)) are robust under any short-range, locally–finite-rank perturbation that does not close the spectral gap (and preserves the symmetry). It follows from the K-theory of (real) Roe C*-algebra (cf. (3.21)) that such a topological phase can be described completely by a $\mathbb{Z}$ or $\mathbb{Z}/2$-valued index.
The groupoid model $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$, on the other hand, has more involved K-theory and numerical invariants. One natural question is in which sense these topological phases are still robust. Moreover, do they lead to interesting invariants that do not occur in the coarse-geometric model? Questions related to robustness can be answered by showing that the numerical index is still continuous, even for some more general perturbations of the Hamiltonian (which does not have to belong to the observable C*-algebra), together with a “quantisation” statement that the range of the numerical index is $\mathbb{Z}$. This implies that such perturbations leave the numerical index unchanged.
We describe a different approach to this robustness question by comparing the groupoid model with the coarse-geometric model on $\Lambda$ using the regular representation. The numerical invariants defined in [Bourne-Mesland:Aperiodic] factor through this representation. Therefore, since $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$ is “robust”, one only needs to understand which topological phases in $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$ still survive there. To this end, we introduce a handy definition of position spectral triples (Definition 2.16). Both the groupoid model and the coarse-geometric model possess such type of spectral triples. In particular, the position spectral triple $\xi^{\mathrm{Roe}}_{\omega}$ over the Roe C*-algebra $\mathrm{C}^{*}_{\mathrm{Roe}}(\omega)$ induces an isomorphism $\operatorname{KK}(\mathrm{C}\ell_{d,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\omega))\to\mathbb{Z}$ (Theorem 3.32). Then we show that those numerical invariants of the groupoid model, given by localising the “bulk cycle” $\prescript{}{d}{\lambda}_{\Omega_{0}}$ at $\omega\in\Omega_{0}$, are pullbacks of the class of $\xi^{\mathrm{Roe}}_{\omega}$ in K-homology under the comparison *-homomorphism $\mathbb{M}_{N}(\mathrm{C}^{*}(\mathcal{G}_{\Lambda}))\to\mathrm{C}^{*}_{\mathrm{Roe}}(\omega)$ (Theorem 4.1). Such invariants are therefore strong in the sense of [Ewert-Meyer:Coarse_geometry].
The comparison *-homomorphisms $\pi_{\omega}^{N}\colon\mathbb{M}_{N}\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\to\mathrm{C}^{*}_{\mathrm{Roe}}(\omega)$ are (localised) regular representation on the tight-binding Hilbert spaces $\ell^{2}(\omega,\mathcal{K})$, where $\mathcal{K}$ is an auxilliary separable Hilbert space, considered as a generalised version of the “fundamental domain”. Such a representation, together with the position operators thereon, are zipped into the data of position spectral triples, which is compatible with the numerical invariants constructed in [Bourne-Mesland:Aperiodic, Ewert-Meyer:Coarse_geometry]. On the other hand, Kubota [Kubota:Controlled_topological_phases] has described a way to describe a comparison *-homomorphism $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\to\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$ whenever $\Lambda$ is aperiodic. We explain in Example 4.4 that Kubota’s construction can be concretely realised as $\pi_{\Lambda}$; in particular, this works for any $\omega\in\Omega_{0}$ disregarding whether or not the corresponding Delone set $\omega$ is aperiodic.
As another application, we show in Theorem 4.11 that certain numerical invariants on Delone sets, coming from “stacking” lower-dimensional topological phases, must be weak, i.e. unstable under perturbation. This generalises a result in [Ewert-Meyer:Coarse_geometry], which explains why certain topological phases of the periodic model ought to be weak. Ewert and Meyer compare the periodic model and the coarse-geometric model using an injective *-homomorphism $\mathrm{C}^{*}(\mathbb{Z}^{d})\hookrightarrow\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d})$. It vanishes on all but one $\mathbb{Z}$-component of K-theory. In particular, topological phases coming from “stacking” lower-dimensional topological phases will all vanish in $\operatorname{K}_{*}(\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d}))$. This allows to conclude that the induced maps $\operatorname{K}_{*}(\mathrm{C}^{*}(\mathbb{Z}^{d}))\to\operatorname{K}_{*}(\mathrm{C}^{*}(\mathbb{Z}^{d+1}))$ also vanishes there. We show that one can replace $\mathbb{Z}^{d+1}$ by a product Delone set of the form $\Lambda\times L$, where $\Lambda\subseteq\mathbb{R}^{m}$ and $L\subseteq\mathbb{R}^{n}$ are both Delone sets. Using the same strategy in [Ewert-Meyer:Coarse_geometry], we show that stacked topological phases factor through the Roe C*-algebra of a flasque space, which has vanishing K-theory. This allows to conclude that such topological phases and their numerical invariants have to vanish in the target Roe C*-algebra.
1.1. Future work
We remark here with some results that will appear in a forthcoming work [Li:Coarse_groupoid_robustness]. Bourne and Mesland [Bourne-Mesland:Aperiodic] have provided another way to construct numerical invariants for the groupoid model of topological phases, by composing the bulk cycle with a semi-finite spectral triple. Unlike the case studied in this present article, such numerical invariants have real numbers as their ranges, and take more than one Delone set $\omega\in\Omega$ as input. We shall show that such numerical invariants are still robust, by using a similar construction of Kubota, but with a family of coarse groupoids.
An analog of Kubota’s construction in the continuum setting is given in [Ewert-Meyer:Coarse_geometry], where the observable C*-algebra $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$ is replaced by $\mathrm{C}^{*}(\Omega_{\Lambda})\rtimes\mathbb{R}^{d}$; and $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$ is replaced by $\mathbb{R}^{d}$. Such constructions are compatible with Kubota’s construction, but one must carefully define the maps between the corresponding “tight-binding” and “continuum” versions of observable C*-algebras.
1.2. Acknowledgement
The author thanks Guo Chuan Thiang, Matthias Ludewig, Matthias Frerichs for helpful discussions and comments. In particular, the author thanks his supervisor Bram Mesland for heartfelt encouragement and a careful proofreading, helping the author to eliminate various errors in an earlier version of this work. The work is part of NWO project 613.009.142 “Noncommutative index theory of discrete dynamical systems”.
1.3. Notation and conventions
We list a few symbols and conventions that will be used in the article.
Dirac notation
Let $\Lambda$ be a discrete set. We write $\lvert x\rangle$ for the function in $\ell^{2}(\Lambda)$ which takes $1$ on $x$ and 00 elsewhere; and $\langle x\rvert$ for the rank-one operator $\ell^{2}(\Lambda)\to\mathbb{C}$ defined by $\lvert y\rangle\mapsto\left\langle x,y\right\rangle$. Let $T\in\mathbb{B}(\ell^{2}(\Lambda))$, then by $\left\langle x\,\middle|T\middle|\,\mathopen{}y\right\rangle$ we shall mean the inner product of $\left\lvert x\right\rangle$ and $T\left\lvert y\right\rangle$. The collection $(T_{x,y})_{x,y\in\Lambda}$ of $T_{x,y}\mathrel{\vcentcolon=}\langle x\,|T|\,\mathopen{}y\rangle$ are called the matrix elements of $T$.
Tensor products
Let $A$ and $B$ be $\mathbb{Z}/2$-graded “real”C*-algebras. We will write $A\otimes B$ for their graded, minimal tensor product. The grading will cause little difference in this article: we will mostly consider $\mathbb{Z}/2$-graded C*-algebras of the form $A\otimes\mathbb{C}\ell_{p,q}$ or $A\otimes\mathrm{C}\ell_{p,q}$, where $A$ is trivially graded. Then the graded tensor product agrees with the ungraded version. We also write $A\odot B$ for the algebraic tensor product of $A$ and $B$.
Groupoids
A groupoid will sometimes be written as $\mathcal{G}\rightrightarrows\mathcal{G}^{0}$ where $\mathcal{G}^{0}$ is its unit space. If $X,Y\subseteq\mathcal{G}^{0}$, then we write $\mathcal{G}_{X}\mathrel{\vcentcolon=}s^{-1}(X)$, $\mathcal{G}^{Y}\mathrel{\vcentcolon=}r^{-1}(Y)$ and $\mathcal{G}_{X}^{Y}\mathrel{\vcentcolon=}\mathcal{G}_{X}\cap\mathcal{G}^{Y}$. We write $\mathrm{C}^{*}(\mathcal{G})$ for the reduced groupoid C*-algebra of $\mathcal{G}$.
Roe C*-algebras
We write $\mathrm{C}^{*}_{\mathrm{Roe}}(X,\mathcal{H}_{X})$ for the Roe C*-algebra defined by an ample module $(X,\mathcal{H}_{X})$. This differs from the usual convention (e.g. in [Willett-Yu:Higher_index_theory]), and agrees with the convention in [Ewert-Meyer:Coarse_geometry]. In particular, we highlight that $\mathrm{C}^{*}(\mathbb{Z}^{d})$ refers to the group C*-algebra of the group $\mathbb{Z}^{d}$, whereas $\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d})$ refers to the Roe C*-algebra of the discrete metric space $\mathbb{Z}^{d}$.
2. Topological phases and their numerical invariants
When we consider physical systems with anti-unitary symmetries, like time-reversal or particle-hole symmetries, then we must represent them as anti-unitary operators on a Hilbert space which commute or anti-commute with the Hamiltonian. This turns the Hilbert space into a “real” Hilbert space, namely, a Hilbert space that carries an involutive, conjugate-linear map. A concrete C*-algebra on this “real” Hilbert space thus becomes a “real” C*-algebra, that is, a C*-algebra together with a involutive, conjugate-linear *-automorphism, called a “real” structure. Kellendonk [Kellendonk:Cst-algebraic_topological_phases] has shown that the topological phases of such a system are classified by a van Daele’s K-theory [Van_Daele:K-theory_1, Van_Daele:K-theory_2], a model of K-theory that works for graded and “real” Banach algebras, of the corresponding observable “real” C*-algebra. We remark that van Daele’s K-theory coincides with Kasparov theory of “real” C*-algebras, cf. [Roe:Pascheke_duality, Bourne-Kellendonk-Rennie:Cayley_transform, Joseph-Meyer:Geometric_van_Daele_K-theory]. Thus we do not have to distinguish that from the “usual” K-theory if we define the latter using Kasparov theory.
The K-theory $\operatorname{KR}_{*}(A,\mathfrak{r})$ of a “real” C*-algebra $(A,\mathfrak{r})$ is by definition the KO-theory $\operatorname{KO}_{*}(A^{\mathfrak{r}})$ of the real C*-algebra $A^{\mathfrak{r}}$, the fixed-point algebra of the “real” structure $\mathfrak{r}$. In our context, all “real” C*-algebras have simple “real” structures, allowing us to focus on their real subalgebras, and we will simply write $\operatorname{K}_{*}(A)$ for their KO-theory.
Similar to complex C*-algebras, real C*-algebras have a bivariant theory, referred to as KKO-theory; we shall write KK for simplicity when the C*-algebra being real is clear from the context. The difference is that it has a different Bott periodicity. That means, taking suspensions or tensoring with Clifford factors on different entries of this bivariant theory usually yield different groups. Thus we will avoid using $\operatorname{KK}_{n}(A,B)$ and only write $\operatorname{KK}(A\otimes\mathrm{C}\ell_{n,0},B)$ or $\operatorname{KK}(A,B\otimes\mathrm{C}\ell_{n,0})$.
This bivariant theory allows us to describe the index pairings in a formal way. Namely, via the Kasparov product
$\operatorname{KK}(\mathrm{C}\ell_{p,q},A)\times\operatorname{KK}(A\otimes\mathrm{C}\ell_{p^{\prime},q^{\prime}},\mathbb{R})\to\operatorname{KK}(\mathrm{C}\ell_{p+p^{\prime},q+q^{\prime}},\mathbb{R}).$
The range can be identified with a certain real K-theory group of $\mathbb{R}$ by real Bott periodicity.
In our case, the first class in $\operatorname{KK}(\mathrm{C}\ell_{p,q},A)$ comes from the physical Hamiltonian $H$. This is a self-adjoint operator affiliated to $A$, which has a spectral gap at $\mu\in\mathbb{R}$. The symmetries of the system are implemented by the Clifford action. The second class is described by a real spectral triple coming from a tight-binding position operator, acting on the tight-binding Hilbert space $\ell^{2}(\Lambda,\mathcal{K})$. In particular, the assumptions for a spectral triple guarantees that any “physically reasonable” observable C*-algebra $A$ must sit inside the Roe C*-algebra $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$, as will be shown in Theorem 3.23.
2.1. “Real” and real C*-algebras
We recall the definition of “real” and real C*-algebras. In other literature, a “real” C*-algebra is also referred to as a Real C*-algebra (cf. [Bourne-Kellendonk-Rennie:Cayley_transform], note the upper case R) or a $\mathrm{C}^{*,\mathrm{r}}$-algebra (cf. [Kellendonk:Cst-algebraic_topological_phases]).
Definition 2.1 ([Kellendonk:Cst-algebraic_topological_phases]*Definition 3.7).
A “real” structure on a $\mathbb{Z}/2$-graded C*-algebra $A$ is a conjugate-linear, grading-preserving *-automorphism $\mathfrak{r}\colon A\to A$ of order 2. A “real” C*-algebra is a C*-algebra together with a “real” structure on it.
When there is no ambiguity, we shall use $A$ alone to denote the “real” C*-algebra $(A,\mathfrak{r})$. For example, if $A=\mathbb{M}_{n}\mathbb{C}$ is the C*-algebra of complex $n\times n$-matrices, then the entrywise complex conjugation defines a “real” structure on $A$.
Definition 2.2.
A “real” involution on a $\mathbb{Z}/2$-graded Hilbert space $\mathcal{H}$ is a conjugate-linear, grading-preserving automorphism of $\mathcal{H}$ of order two.
Let $\Theta$ be a “real” involution on a $\mathbb{Z}/2$-graded Hilbert space. Then $\mathbb{B}(\mathcal{H})$ becomes a $\mathbb{Z}/2$-graded “real” C*-algebra with “real” structure $T\mapsto\Theta T\Theta$.
Definition 2.3.
A representation of a $\mathbb{Z}/2$-graded “real” C*-algebra $A$ is a *-homomorphism $\pi\colon A\to\mathbb{B}(\mathcal{H})$ for a $\mathbb{Z}/2$-graded “real” Hilbert space $\mathcal{H}$, which intertwines both the $\mathbb{Z}/2$-gradings and the “real” structures.
Now we define real C*-algebras and real Hilbert spaces.
Definition 2.4.
A real Hilbert space is a Hilbert space over $\mathbb{R}$. A real C*-algebra is a norm-closed subalgebra of $\mathbb{B}(\mathcal{H_{\mathbb{R}}})$, where $\mathcal{H_{\mathbb{R}}}$ is a Hilbert space over $\mathbb{R}$.
If $A$ is a “real” C*-algebra, then the collection of its real elements
$A^{\mathfrak{r}}\mathrel{\vcentcolon=}\left\{a\in A\;\middle|\;\mathfrak{r}(a)=a\right\}$
is a real C*-algebra. Conversely, if $A_{\mathbb{R}}$ is a real C*-algebra, then $A_{\mathbb{R}}\otimes_{\mathbb{R}}\mathbb{C}$ is a “real” C*-algebra with “real” structure
$\mathfrak{r}(a\otimes_{\mathbb{R}}z)\mathrel{\vcentcolon=}a\otimes_{\mathbb{R}}\overline{z}.$
Let $(X,\tau)$ be an locally compact, Hausdorff involutive space, that is, a locally compact Hausdorff space $X$ together with a homeomorphism $\tau\colon X\to X$ satisfying $\tau^{2}=\operatorname{id}_{X}$. If $X$ is a manifold, then we also say $(X,\tau)$ is a “real” manifold. The C*-algebra $\mathrm{C}_{0}(X)$ carries a “real” structure
$\mathfrak{r}(f)(x)\mathrel{\vcentcolon=}\overline{f(\tau(x))},$
yielding a “real” C*-algebra $(\mathrm{C}_{0}(X),\mathfrak{r})$ as well as its corresponding real C*-algebra $\mathrm{C}_{0}(X)^{\mathfrak{r}}$. The “real” Gelfand–Naimark theorem due to Arens and Kaplansky [Arens_Kaplansky:Topological_representation]*Theorem 9.1 asserts that every commutative “real” C*-algebra is isomorphic to $(\mathrm{C}_{0}(X),\mathfrak{r})$ for some involutive space $(X,\tau)$; and every commutative real C*-algebra is isomorphic to $\mathrm{C}_{0}(X)^{\mathfrak{r}}$.
2.2. Clifford algebras
Let $\mathbb{C}\ell_{p,q}$ be the finite-dimensional graded “real” C*-algebra generated by $\gamma^{1},\dots,\gamma^{p},\rho^{1},\dots,\rho^{q}$, satisfying:
  • $\gamma^{1},\dots,\gamma^{p}$ are odd, self-adjoint, involutive and real;
  • $\rho^{1},\dots,\rho^{q}$ are odd, anti-self-adjoint, anti-involutive and real;
  • $\gamma^{1},\dots,\gamma^{p},\dots,\rho^{1},\dots,\rho^{q}$ mutually anti-commute.
That is, we require that for all $i,j$:
$(\gamma^{j})^{2}=1,\quad$ $(\gamma^{j})^{*}=\gamma^{j},\quad$ $\mathfrak{r}(\gamma^{j})=\gamma^{j};$
$(\rho^{j})^{2}=-1,\quad$ $(\rho^{j})^{*}=-\rho^{j},\quad$ $\mathfrak{r}(\rho^{j})=\rho^{j}.$
The real subalgebra of $\mathbb{C}\ell_{p,q}$ is the $\mathbb{R}$-algebra generated by the same generators and relations. We write $\mathrm{C}\ell_{p,q}\mathrel{\vcentcolon=}(\mathbb{C}\ell_{p,q})^{\mathfrak{r}}$ for this $\mathbb{Z}/2$-graded, real C*-algebra. Up to Morita equivalence of graded real C*-algebras, there are only eight possible $\mathrm{C}\ell_{p,q}$ satisfying
$\mathrm{C}\ell_{p,q}\otimes\mathrm{C}\ell_{p^{\prime},q^{\prime}}\simeq\mathrm{C}\ell_{p+p^{\prime},q+q^{\prime}}.$
Moreover, $\mathrm{C}\ell_{1,1}$ is isomorphic to the $\mathbb{Z}/2$-graded real C*-algebra $\mathbb{M}_{2}(\mathbb{R})$, whose grading is given by diagonal–off-diagonal elements. Thus
$\mathrm{C}\ell_{d,d}\simeq\mathbb{M}_{2}(\mathbb{R})\otimes\mathbb{M}_{2}(\mathbb{R})\otimes\dots\otimes\mathbb{M}_{2}(\mathbb{R})\simeq\mathbb{M}_{2^{d}}(\mathbb{R}).$
Kasparov has constructed a canonical representation of $\mathbb{C}\ell_{p,q}$ in [Kasparov:K-functor_extension]. Let $\mathbb{C}^{d}$ be the “real” Hilbert space with basis $e_{1},\dots,e_{d}$ and “real” involution
$\sum_{i=1}^{d}c_{i}e_{i}\;\longmapsto\;\sum_{i=1}^{d}\overline{c_{i}}e_{i}.$
Let $\bigwedge\nolimits^{*}\mathbb{C}^{d}$ be the exterior algebra of $\mathbb{C}^{d}$. It is graded by the subspace of odd or even differential forms $\bigwedge\nolimits^{*}\mathbb{C}^{d}=\bigwedge\nolimits^{\text{odd}}\mathbb{C}^{d}\oplus\bigwedge\nolimits^{\text{even}}\mathbb{C}^{d}$. The “real” structure on $\mathbb{C}^{d}$ extends to $\bigwedge\nolimits^{*}\mathbb{C}^{d}$, that is,
$\sum_{i_{i}i_{2}\dots i_{k}}a_{i_{1}i_{2}\dots i_{k}}e_{i_{1}}\wedge e_{i_{2}}\wedge\dots\wedge e_{i_{k}}\longmapsto\sum_{i_{i}i_{2}\dots i_{k}}\overline{a_{i_{1}i_{2}\dots i_{k}}}e_{i_{1}}\wedge e_{i_{2}}\wedge\dots\wedge e_{i_{k}},$
turning $\bigwedge\nolimits^{*}\mathbb{C}^{d}$ into a $\mathbb{Z}/2$-graded “real” Hilbert space.
Definition and Lemma 2.5.
Let $j\in\{1,\dots,d\}$ and $\lambda_{j}\colon\bigwedge\nolimits^{*}\mathbb{C}^{d}\to\bigwedge\nolimits^{*}\mathbb{C}^{d}$ be the exterior product with $e_{j}$, that is, $\lambda_{j}(\omega)=e_{j}\wedge\omega$. Then its adjoint $\lambda_{j}^{*}\colon\bigwedge\nolimits^{*}\mathbb{C}^{d}\to\bigwedge\nolimits^{*}\mathbb{C}^{d}$ is contraction with $e_{j}$, that is, $\lambda_{j}^{*}(\omega)=e_{j}\mathbin{\lrcorner}\omega$.
As a consequence, there is a representation of the $\mathbb{Z}/2$-graded “real” C*-algebra $\mathbb{C}\ell_{p,q}$ on $\bigwedge\nolimits^{*}\mathbb{C}^{p+q}$, sending $\gamma^{j}$ to $\lambda_{j}+\lambda_{j}^{*}$ and $\rho^{j}$ to $\lambda_{p+j}-\lambda_{p+j}^{*}$. We call it the standard representation of $\mathbb{C}\ell_{p,q}$.
Proof.
The operators $\lambda_{j}$ and their adjoints satisfy
$\lambda_{i}\lambda_{j}+\lambda_{j}\lambda_{i}=0,\quad\lambda_{i}^{*}\lambda_{j}+\lambda_{j}\lambda_{i}^{*}=\left\langle e_{i},e_{j}\right\rangle.$
Moreover, $\lambda_{j}$ and $\lambda_{j}^{*}$ are odd and real for all $j$. So we have
$(\lambda_{j}+\lambda_{j}^{*})^{2}=1,\quad(\lambda_{j}-\lambda_{j}^{*})^{2}=-1,\quad(\lambda_{i}+\lambda_{i}^{*})(\lambda_{p+j}-\lambda_{p+j}^{*})=0.$
Thus the operators $\lambda_{1}+\lambda_{1}^{*},\dots\lambda_{r}+\lambda_{r}^{*},\dots,\lambda_{p+1}-\lambda_{p+1}^{*},\dots\lambda_{p+q}-\lambda_{p+q}^{*}$ generate a copy of $\mathbb{C}\ell_{p+q}$ in $\mathbb{B}(\bigwedge\nolimits^{*}\mathbb{C}^{p+q})$. ∎
Next, we describe the representation of the graded “real” C*-algebra $\mathrm{C}^{*}(\mathbb{Z}^{d})\otimes\mathbb{C}\ell_{0,d}$ following [Ewert-Meyer:Coarse_geometry]*Section 4. The grading on $\mathrm{C}^{*}(\mathbb{Z}^{d})\otimes\mathbb{C}\ell_{0,d}$ is the tensor product of the trivial grading on $\mathrm{C}^{*}(\mathbb{Z}^{d})$ with the standard grading on $\mathbb{C}\ell_{0,d}$. The “real” structure on $\mathrm{C}^{*}(\mathbb{Z}^{d})$ is given by the pointwise complex conjugation, that is,
$\mathfrak{r}(f)(n)\mathrel{\vcentcolon=}\overline{f(n)},\quad n\in\mathbb{Z}^{d}.$
Then the regular representation of the complex C*-algebra $\mathrm{C}^{*}(\mathbb{Z}^{d})$ extends to a *-representation on the “real” Hilbert space $\ell^{2}(\mathbb{Z}^{d})$, whose “real” structure is given by pointwise conjugation. The “real” structure on $\mathrm{C}^{*}(\mathbb{Z}^{d})\otimes\mathbb{C}\ell_{0,d}$ is the tensor product of this “real” structure with the standard one on $\mathbb{C}\ell_{0,d}$.
The Fourier transform maps the “real” C*-algebra $\mathrm{C}^{*}(\mathbb{Z}^{d})$ to the “real” C*-algebra $\mathrm{C}(\mathbb{T}^{d})$. Here $\mathbb{T}^{d}$ is a “real” manifold, with involution $z\mapsto\overline{z}$.
Lemma 2.6 (cf. [Ewert-Meyer:Coarse_geometry]*Section 4).
There is an isomorphism of $\mathbb{Z}/2$-graded, “real” Hilbert spaces
(2.7) $U\colon\ell^{2}(\mathbb{Z}^{d})\otimes\bigwedge\nolimits^{*}\mathbb{C}^{d}\xrightarrow{\sim}L^{2}\left(\bigwedge\nolimits^{*}\mathrm{T}^{*}\mathbb{T}^{d}\right)$
given by
$\left\lvert k\right\rangle\otimes e_{i_{1}}\wedge\dots e_{i_{l}}\mapsto\frac{z^{k}}{z_{i_{1}}\dots z_{i_{l}}}\mathrm{d}z_{i_{1}}\wedge\dots\wedge\mathrm{d}z_{i_{l}},$
where
$z^{k}\mathrel{\vcentcolon=}z_{1}^{k_{1}}z_{2}^{k_{2}}\dots z_{d}^{k_{d}},\quad k=(k_{1},\dots,k_{d})\in\mathbb{Z}^{d}.$
This induces an isomorphism of $\mathbb{Z}/2$-graded “real” C*-algebras
$\mathrm{C}^{*}(\mathbb{Z}^{d})\otimes\mathbb{C}\ell_{0,d}\simeq\mathrm{C}(\mathbb{T}^{d})\otimes\mathbb{C}\ell_{0,d}\simeq\mathrm{C}(\mathbb{T},\mathbb{C}\ell(\mathbb{T}^{d}))$
where $\mathbb{C}\ell(\mathbb{T}^{d})$ is the “real” Clifford bundle of the “real” manifold $\mathbb{T}^{d}$. In particular, under this isomorphism, the canonical representation of $\mathbb{C}\ell_{0,d}$ as in 2.5 is mapped to the Clifford multiplication of $\mathbb{C}\ell(\mathbb{T}^{d})$ on $L^{2}\left(\bigwedge\nolimits^{*}\mathrm{T}^{*}\mathbb{T}^{d}\right)$.
Let ${\scriptstyle\mathsf{X}}_{j}$ be the $j$-th self-adjoint position operator on $\ell^{2}(\mathbb{Z}^{d})$ defined by
${\scriptstyle\mathsf{X}}_{j}f(k)\mathrel{\vcentcolon=}k_{j}f(k)$
with domain the space of rapidly decaying functions $\mathbb{Z}^{d}\to\mathbb{C}$. Thus the operator ${\scriptstyle\mathsf{X}}=\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}$ is odd, self-adjoint acting on the $\mathbb{Z}/2$-graded “real” Hilbert space $\ell^{2}(\mathbb{Z}^{d})\otimes\bigwedge\nolimits^{*}\mathbb{C}^{d}$.
Lemma 2.8.
The Fourier transform (2.7) induces a unitary equivalence between the unbounded operator ${\scriptstyle\mathsf{X}}={\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}$ on $\ell^{2}(\mathbb{Z}^{d})\otimes\bigwedge^{*}\mathbb{C}^{d}$ and the Hodge–de Rham operator $\mathrm{d}+\mathrm{d}^{*}$ on $L^{2}(\bigwedge^{*}\mathrm{T}^{*}\mathbb{T}^{d})$. Both operators preserve the real subspace $\ell^{2}(\mathbb{Z}^{d})_{\mathbb{R}}\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d}\simeq L^{2}(\bigwedge^{*}\mathrm{T}^{*}\mathbb{T}^{d})_{\mathbb{R}}$, and hence are “real”.
Proof.
By definition, we have
${\scriptstyle\mathsf{X}}_{j}\otimes\lambda_{j}\left(\left\lvert k\right\rangle\otimes e_{i_{1}}\wedge\dots\wedge e_{i_{l}}\right)$ $=k_{j}\left\lvert k\right\rangle\otimes e_{j}\wedge e_{i_{1}}\wedge\dots\wedge e_{i_{l}}$
$=U^{*}\left(k_{j}z^{k}\cdot\frac{\mathrm{d}z_{j}}{z_{j}}\wedge\frac{\mathrm{d}z_{i_{1}}}{z_{i_{1}}}\wedge\dots\wedge\frac{\mathrm{d}z_{i_{l}}}{z_{i_{l}}}\right).$
The de Rham operator $\mathrm{d}$ satisfies
$\mathrm{d}\left(\frac{z^{k}}{z_{i_{1}}\dots z_{i_{l}}}\mathrm{d}z_{i_{1}}\wedge\dots\wedge\mathrm{d}z_{i_{l}}\right)$ $=\sum_{j=1}^{d}z_{1}^{k_{1}}\cdots(k_{j}z_{j}^{k_{j}-1}\mathrm{d}z_{j})\cdots z_{d}^{k_{d}}\wedge\frac{\mathrm{d}z_{i_{1}}}{z_{i_{1}}}\wedge\dots\wedge\frac{\mathrm{d}z_{i_{l}}}{z_{i_{l}}}.$
$=\sum_{j=1}^{d}k_{j}z^{k}\cdot\frac{\mathrm{d}z_{j}}{z_{j}}\wedge\frac{\mathrm{d}z_{i_{1}}}{z_{i_{1}}}\wedge\dots\wedge\frac{\mathrm{d}z_{i_{l}}}{z_{i_{l}}}.$
So $U^{*}\mathrm{d}U$ acts by $\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\lambda_{j}$ on $\ell^{2}(\mathbb{Z}^{d})\otimes\bigwedge\nolimits^{*}\mathbb{C}^{d}$. Therefore, $\mathrm{d}+\mathrm{d}^{*}$ is unitarily equivalent to the operator
$\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes(\lambda_{j}+\lambda_{j}^{*})=\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}$
by the standard representation of $\mathbb{C}\ell_{d,0}$ in 2.5. The generators $\gamma^{j}$ are “real” in $\mathbb{C}\ell_{d,0}$, hence maps $\bigwedge\nolimits^{*}\mathbb{R}^{d}$ to $\bigwedge\nolimits^{*}\mathbb{R}^{d}$. The $j$-th position operator ${\scriptstyle\mathsf{X}}_{j}$ preserves $\ell^{2}(\mathbb{Z}^{d})_{\mathbb{R}}$ because ${\scriptstyle\mathsf{X}}_{j}\left\lvert k\right\rangle=k_{j}\left\lvert k\right\rangle$ and $k_{j}$ is real for any $k\in\mathbb{Z}^{d}$. Therefore the operator ${\scriptstyle\mathsf{X}}=\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}$ preserves $\ell^{2}(\mathbb{Z}^{d})_{\mathbb{R}}\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d}$. ∎
2.3. Topological phases and K-theory
We briefly recall the noncommutative framework for symmetry-protected topological phases of matter. Let $\mathcal{H}$ be a complex Hilbert space, which we regard as the “physically relevant” one. The $\mathsf{CT}$-symmetry of a physical system is described by:
  • a finite group $G\subseteq\mathbb{Z}/2\times\mathbb{Z}/2$;
  • a group homomorphism $\varphi\colon G\to\{\pm 1\}$;
  • a 2-cocycle $\sigma\colon G\times G\to\mathbb{T}$;
  • a Borel map $g\mapsto\theta_{g}$ from $G$ to the set of maps on $\mathcal{H}$
such that
  • $\theta_{g}$ is a unitary (resp. anti-unitary) operator on $\mathcal{H}$ if $\varphi(g)=+1$ (resp.$\varphi(g)=-1$ );
  • $\theta_{g_{1}}\theta_{g_{2}}=\sigma(g_{1},g_{2})\theta_{g_{1}g_{2}}$.
The tuple $(G,\varphi,\sigma,\theta)$ is called a projective unitary/anti-unitary (PUA) representation of the $\mathsf{CT}$-symmetry group $G$ , cf. [Thiang:K-theoretic_classification, Bourne-Carey-Rennie:NC_framework_topological_insulators].
Let $A$ be a $\mathbb{Z}/2$-graded “real” C*-subalgebra of $\mathbb{B}(\mathcal{H})$ An abstract insulator in $A$ is a self-adjoint element $h\in A$ which has a spectral gap at $\mu$, and such that $h$ is compatible with the $\mathsf{CT}$-symmetry group. The characteristic function of $(-\infty,\mu)$ is continuous on the spectrum of $h$, thus continuous functional calculus gives an element $p_{\mu}\mathrel{\vcentcolon=}\chi_{(-\infty,\mu)}(h)$, called the Fermi projection. The following proposition comes from [Kellendonk:Cst-algebraic_topological_phases]*Section 6.
Proposition 2.9.
Let $h$ be a self-adjoint element in a $\mathbb{Z}/2$-graded “real” C*-algebra $(A,\mathfrak{r})$, which has a spectral gap at $\mu$. If $h$ has no symmetry, then $h-\mu$ represents a class in $\operatorname{K}_{0}(A)$, given by its Fermi projection $p_{\mu}$; if $h$ has a chiral symmetry, then $h-\mu$ represents a class in $\operatorname{K}_{1}(A)$; if $h$ has anti-unitary symmetries, then $h-\mu$ represents a class in $\operatorname{K}_{n}(A^{\mathfrak{r}})$, the degree $n$ depends on the $\mathsf{CT}$-symmetry group of the system.
In order to compute the index pairing in a compatible way with [Bourne-Mesland:Aperiodic]*Section 6.2, we use real Kasparov modules to represent real K-theory classes, as in [Bourne-Carey-Rennie:NC_framework_topological_insulators]*Section 3.3. In particular, the real Kasparov modules in this approach consist of a finitely generated projective $A$-module, and such that the operator of this Kasparov module is zero. We remark here that the classification of $\mathsf{CT}$-symmetry groups in the setting of [Bourne-Carey-Rennie:NC_framework_topological_insulators] is different from Kellendonk, as discussed in [Kellendonk:Cst-algebraic_topological_phases]*Section 1.1.1, as it will change the degree shift in the K-theory class of Hamiltonians that have particle-hole symmetries. Since we are not concerned with the precise degree in K-theory in this article, this subtle difference will not affect our results.
[…]
Let $h$ be an abstract insulator, compatible with the $\mathsf{CT}$-symmetry group $G$. Then there exist:
  • a projection $p\in\mathbb{M}_{N}(A)$;
  • a natural number $n$ and $N\in\{1,2,4\}$, together with a representation
    $\mathrm{C}\ell_{n,0}\to\mathbb{B}(pA^{\oplus N}),$
    which are determined by the $\mathsf{CT}$-symmetry group $G$ as in [Bourne-Carey-Rennie:NC_framework_topological_insulators]*table 1;
such that the following finitely generated Kasparov module represents a class $[h]\in\operatorname{KK}(\mathrm{C}\ell_{d,0},A)$:
[…]
If the system contains at most time-reversal symmetry, that is, $G\subseteq\{1,\mathsf{T}\}$, then the projection $p=p_{\mu}$ is the Fermi projection of $h$. If the system carries particle-hole symmetry or chiral symmetry, then the projection $p$ is closely related, but not equal, to $p_{\mu}$.
2.4. Position spectral triples and index pairings
In order to extact a numerical invariant from a given topological phase, described by the class $[h]\in\operatorname{KK}(\mathrm{C}\ell_{n,0},A)$ of an abstract insulator, one seeks a map
$\operatorname{KK}(\mathrm{C}\ell_{n,0},A)\to\mathbb{R}.$
One way to construct such a map is by pairing with a suitable K-homology class, represented by a (real) spectral triple:
Definition 2.12.
Let $A$ be a $\mathbb{Z}/2$-graded real C*-algebra. A (real) spectral triple $(\mathcal{A},\mathcal{H},D)$ over $A$ consists of:
  • A $\mathbb{Z}/2$-graded real Hilbert space $\mathcal{H}$, together with a grading-preserving *-representation $\varphi\colon A\to\mathbb{B}(\mathcal{H})$;
  • A dense *-subalgebra $\mathcal{A}\subseteq A$;
  • An unbounded, self-adjoint odd operator $D\colon\operatorname{dom}D\subseteq\mathcal{H}\to\mathcal{H}$
such that:
  • $\varphi(a)(1+D^{2})^{-1}$ is compact for all $a\in\mathcal{A}$;
  • For every $a\in\mathcal{A}$, $\varphi(a)$ maps $\operatorname{dom}D$ into $\operatorname{dom}D$; and $[D,\varphi(a)]$ extends to a bounded operator on $\mathcal{H}$.
Given $p>0$, we say a spectral triple $(\mathcal{A},\mathcal{H},D)$ is $p$-summable, if $(1+D^{2})^{-s/2}$ is trace class for all $s>p$.
A real spectral triple represents a class in real K-homology. More precisely, Kasparov has already defined his bivariant theory for “real” C*-algebras in [Kasparov:K-functor_extension]. Namely, given two $\mathbb{Z}/2$-graded “real” C*-algebras $A$ and $B$, Kasparov defined the KKR-group $\operatorname{KKR}(A,B)$, whose elements are equivalence class of “real” Kasparov $A$-$B$–modules. Restricting to the real elements $A_{\mathbb{R}}$ and $B_{\mathbb{R}}$ of $A$ and $B$ yields a real Kasparov module of $\operatorname{KKO}(A_{\mathbb{R}},B_{\mathbb{R}})$, which has the same definition as a (complex) Kasparov module, but with all Hilbert spaces and C*-algebras being real. As explained in [Bourne-Kellendonk-Rennie:Cayley_transform]*Definition 2.8, restricting to real parts does not lose any information of a “real” Kasparov module, and $\operatorname{KKR}(A,B)\simeq\operatorname{KKO}(A_{\mathbb{R}},B_{\mathbb{R}})$.
A real spectral triple over $A$ is the same thing as a real unbounded Kasparov $(A,\mathbb{R})$-module. An unbounded (real) Kasparov module is, via the bounded transform, mapped to a Fredholm module, which represents a class in K-homology. The Kasparov product between a K-theory class and a K-homology class is also referred to as an index pairing. The range is the K-theory of $\mathbb{R}$ with a degree shift, which can equal $\mathbb{Z}$, $\mathbb{Z}/2$ or 00. Thus a real spectral triple for a real C*-algebra $A$ generates a number from a topological phase in $A$ via the Kasparov product.
We note that the Kasparov product is functorial in the following sense:
Lemma 2.13.
Let $f\colon A\to B$ be a *-homomorphism between real C*-algebras. Let $x$ be a class in $\operatorname{KK}(B\otimes\mathrm{C}\ell_{0,d},\mathbb{R})$. Let $f^{*}x$ be the Kasparov product of $f\in\operatorname{KK}(A,B)$ and $x\in\operatorname{KK}(B\otimes\mathrm{C}\ell_{0,d},\mathbb{R})$. Then the following diagram commutes:
${\operatorname{KK}(\mathrm{C}\ell_{d,0},A)}$${\operatorname{KK}(\mathrm{C}\ell_{d,d},\mathbb{R})}$${\operatorname{KK}(\mathrm{C}\ell_{d,0},B)}$$\scriptstyle{f^{*}x}$$\scriptstyle{f}$$\scriptstyle{x}$
where all arrows are given by taking Kasparov products.
We call the class $f^{*}x$ the pullback of $x\in\operatorname{KK}(B\otimes\mathrm{C}\ell_{0,d},\mathbb{R})$ along $f$. When $x$ is given by a real spectral triple $(\mathcal{B},\mathcal{H},D)$, and $B$ is represented on $\mathcal{H}$ by $\varphi\colon B\to\mathbb{B}(\mathcal{H})$, then the *-algebra $\mathcal{A}=f^{-1}(\mathcal{B})$ is dense in $\mathcal{A}$, and the class of $f^{*}x$ is represented by the spectral triple $(\mathcal{A},\mathcal{H},D)$ where $A$ is represented on $\mathcal{H}$ by $\varphi\circ f\colon A\to\mathbb{B}(\mathcal{H})$.
In the following, we shall look at a class of spectral triples, which we call (tight-binding) position spectral triples as they are built from the position operators on the corresponding Hilbert space of a Delone set $\Lambda$.
Definition 2.14.
A discrete infinite set $\Lambda\subseteq\mathbb{R}^{d}$ is called a Delone set, if there exists positive real numbers $R$ and $r$, such that
#$\#(\operatorname{B}(x,r)\cap\Lambda)\leq 1,\quad\#(\operatorname{B}(x,R)\cap\Lambda)\geq 1$
for all $x\in\mathbb{R}^{d}$, where $\operatorname{B}(x,r)$ denotes the open $r$-ball centered at $x$. We also say the set $\Lambda$ is $(r,R)$-Delone.
Denote the collection of $(r,R)$-Delone sets in $\mathbb{R}^{d}$ by $\operatorname{Del}_{(r,R)}(\mathbb{R}^{d})$.
Given a Delone set $\Lambda$, equipped with the subspace metric $d(x,y)\mathrel{\vcentcolon=}\left\lvert x-y\right\rvert$. The Voronoi tiling decomposes $\mathbb{R}^{d}$ into tiles lablled by $x\in\Lambda$, and thus decomposes the real Hilbert space $L^{2}(\mathbb{R}^{d})_{\mathbb{R}}$ into a direct sum:
(2.15) $L^{2}(\mathbb{R}^{d})_{\mathbb{R}}\simeq\bigoplus_{x\in\Lambda}L^{2}(T_{x})_{\mathbb{R}}\simeq\ell^{2}(\Lambda,\mathcal{K})_{\mathbb{R}},$
where $T_{x}$ is (interior of) the tile associated to the point $x\in\Lambda$. If $\Lambda$ is a periodic lattice, then all $T_{x}$’s are translated copies of the fundamental domain of $\Lambda$, where $\Lambda$ is identified with the discrete, cocompact subgroup of $\mathbb{R}^{d}$ generated by translations by vectors $x\in\Lambda$. In this sense, we may regard $T_{x}$’s as a generalised version of “fundamental domains” of the dynamics on the aperiodic lattice $\Lambda$.
We may identify all $L^{2}(T_{x})_{\mathbb{R}}$ with a fixed, separable real Hilbert space $\mathcal{K}$. Thus an operator $T\in\mathbb{B}(L^{2}(\mathbb{R}^{d})_{\mathbb{R}})\simeq\mathbb{B}(\ell^{2}(\Lambda,\mathcal{K})_{\mathbb{R}})$ can be described by its matrix elements $(T_{x,y})_{x,y\in\Lambda}$, where
$T_{x,y}\mathrel{\vcentcolon=}\left\langle x\,\middle|T\middle|\,\mathopen{}y\right\rangle\in\mathbb{B}(\mathcal{K}).$
Definition 2.16.
Let $\Lambda\subseteq\mathbb{R}^{d}$ be a countable discrete subset (e.g. a Delone set). A position spectral triple associated to $\Lambda$ is a real spectral triple of the form
(2.17) $\left(\mathcal{A}\otimes\mathrm{C}\ell_{0,d},\quad\ell^{2}(\Lambda,\mathcal{K})_{\mathbb{R}}\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d},\quad\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}\right),$
where:
  • $A$ is a real C*-algebra, which carries a *-representation $\varphi\colon A\to\mathbb{B}(\ell^{2}(\Lambda,\mathcal{K})_{\mathbb{R}})$ for some separable real Hilbert space $\mathcal{K}$; $\mathcal{A}\subseteq A$ is a dense *-subalgebra;
  • $A\otimes\mathrm{C}\ell_{0,d}$ is represented on $\ell^{2}(\Lambda,\mathcal{K})_{\mathbb{R}}\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d}$ through the tensor product of $\varphi$ and the standard representation of $\mathrm{C}\ell_{0,d}$;
  • ${\scriptstyle\mathsf{X}}_{j}$ is the $j$-th tight-binding “position” operator on $\ell^{2}(\Lambda,\mathcal{K})_{\mathbb{R}}$, given by
    $({\scriptstyle\mathsf{X}}_{j}f)(x)\mathrel{\vcentcolon=}x_{j}f(x),\qquad x=(x_{1},\dots,x_{d})\in\Lambda\subseteq\mathbb{R}^{d};$
  • $\gamma_{1},\dots,\gamma_{d}$ are the generators of $\mathrm{C}\ell_{d,0}$, represented on $\bigwedge\nolimits^{*}\mathbb{R}^{d}$ via the standard representation.
We say a position spectral triple is locally compact, if the matrix elements $\varphi(a)_{x,y}\mathrel{\vcentcolon=}\left\langle x\,\middle|\varphi(a)\middle|\,\mathopen{}y\right\rangle$ are compact for all $a\in\mathcal{A}$ and $x,y\in\Lambda$.
Let $A$ be a real C*-algebra. A position spectral triple of the form (2.17) gives a class in $\operatorname{KK}(A\otimes\mathrm{C}\ell_{0,d},\mathbb{R})$. Via the Kasparov product, it gives a map
$\operatorname{KK}(\mathrm{C}\ell_{d,0},A)\to\operatorname{KK}(\mathrm{C}\ell_{d,d},\mathbb{R})\simeq\mathbb{Z}.$
Example 2.18.
Let $\mathbb{R}[\mathbb{Z}^{d}]$ be the real convolution algebra of $\mathbb{Z}^{d}$. That is, the elements of $\mathbb{R}[\mathbb{Z}^{d}]$ are real-valued, finitely supported functions $f\colon\mathbb{Z}^{d}\to\mathbb{R}$. It carries the following convolution product and involution:
$(f*g)(x)\mathrel{\vcentcolon=}\sum_{x_{1}+x_{2}=x}f(x_{1})g(x_{2}),\quad f^{*}(x)\mathrel{\vcentcolon=}f(-x).$
The left multiplication of $\mathbb{R}[\mathbb{Z}^{d}]$ on itself extends to a injective *-representation
$\lambda\colon\mathbb{R}[\mathbb{Z}^{d}]\to\mathbb{B}(\ell^{2}(\mathbb{Z}^{d})_{\mathbb{R}}).$
The closure of $\lambda(\mathbb{R}[\mathbb{Z}^{d}])$ in the real C*-algebra $\mathbb{B}(\ell^{2}(\mathbb{Z}^{d})_{\mathbb{R}})$ is the real group C*-algebra $\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}}$.
Let $\mathcal{K}$ be a separable real Hilbert space. Fix a natural number $N$ and represent $\mathbb{M}_{N}(\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}})$ as well as its dense subalgebra $\mathbb{M}_{N}(\mathbb{R}[\mathbb{Z}^{d}])$ on $\ell^{2}(\mathbb{Z}^{d},\mathcal{K})_{\mathbb{R}}$ using $e_{N}\circ\lambda^{N}$, where $\lambda^{N}$ is the entrywise extension of $\lambda$ to the matrix algebra $\mathbb{M}_{N}\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}}=\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}}\otimes\mathbb{M}_{N}(\mathbb{R})$; and $e_{N}\colon\mathbb{M}_{N}(\mathbb{R})\to\mathbb{K}(\mathcal{K})$ is any rank-$N$ corner embedding. Then $\langle f,g\rangle\in\mathbb{M}_{N}(\mathbb{R})$ for all $f,g\in\mathbb{R}[\mathbb{Z}]$. It follows that
$\left(\mathbb{M}_{N}(\mathbb{R}[\mathbb{Z}])\otimes\mathrm{C}\ell_{0,d},\quad\ell^{2}(\mathbb{Z}^{d},\mathcal{K})_{\mathbb{R}}\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d},\quad\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}\right)$
is a locally compact position spectral triple associated to the Delone set $\Lambda$. It gives a map
$\operatorname{KK}(\mathrm{C}\ell_{d,0},\mathbb{M}_{N}\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}})\to\operatorname{KK}(\mathrm{C}\ell_{d,d},\mathbb{R})\simeq\mathbb{Z}.$
This spectral triple is used later in Lemma 3.26 from [Ewert-Meyer:Coarse_geometry]*Section 4. It represents the Dirac element of the “real” manifold $\mathbb{T}^{d}$ in the sense of Kasparov [Kasparov:Equivariant_KK-theory]*Definition–Lemma 4.2.
Remark 2.19.
If $\Lambda\subseteq\mathbb{R}^{d}$ is a discrete set but not uniformly discrete, then the Voronoi tiling associated to $\Lambda$ may lack a good physical interpretation. Nevertheless, the above definition still makes sense. This happens, e.g. if we consider the union of two Delone sets $\Lambda_{1}$ and $\Lambda_{2}$, then a position spectral triple associated to $\Lambda\mathrel{\vcentcolon=}\Lambda_{1}\cup\Lambda_{2}$ should be represented on the real Hilbert space $\ell^{2}(\Lambda_{1}\cup\Lambda_{2},\mathcal{K})_{\mathbb{R}}$.
3. Models of aperiodic topological insulators
We describe two observable C*-algebras, modelling a topological insulator that lives on an aperiodic lattice. We assume that the aperiodic lattice is described by a Delone set $\Lambda\subseteq\mathbb{R}^{d}$ (Definition 2.14).
From now on, all Hilbert spaces and C*-algebras are assumed to be real, and we omit their footnotes $\mathbb{R}$ for simplicity.
3.1. Groupoid C*-algebra of a Delone set
In the groupoid approach, a Delone set $\Lambda$ is considered as a point in the space of infinite, discrete subsets in $\mathbb{R}^{d}$. Its dynamics is given by translation of the point set by vectors in $\mathbb{R}^{d}$. More precisely, we identify a Delone set $\Lambda$ with its corresponding pure point measure $\sum_{x\in\Lambda}\delta_{x}$ on $\mathbb{R}^{d}$, thus identify the space of all $(r,R)$-Delone sets as a subspace of $\mathcal{M}(\mathbb{R}^{d})=\mathrm{C}_{\mathrm{c}}(\mathbb{R}^{d})^{\prime}$, the space of all measures on $\mathbb{R}^{d}$, equipped with the weak*-topology. Following [Bellissard-Herrmann-Zarrouati:Hulls_aperiodic_solids], we write:
  • $\mathcal{M}(\mathbb{R}^{d})$ for the space of all measures on $\mathbb{R}^{d}$;
  • $\operatorname{QD}(\mathbb{R}^{d})$ for the set of all pure point measures $\nu$ on $\mathbb{R}^{d}$ such that $\nu(\{x\})\in\mathbb{N}$ for all $x\in\mathbb{R}^{d}$;
  • $\operatorname{UD}_{r}(\mathbb{R}^{d})$ for the subset of $\operatorname{QD}(\mathbb{R}^{d})$ such that $\nu(\operatorname{B}(x;r))\leq 1$ for all $x\in\mathbb{R}^{d}$.
In the context below, we shall not distinguish between a discrete set and its corresponding pure point measure.
Proposition 3.1 ([Bellissard-Herrmann-Zarrouati:Hulls_aperiodic_solids]*Theorem 1.5, Section 2.1).
Using the notation above and fix $0<r<R$, we have the following:
  1. (1)
    there are inclusions of closed sets
    $\operatorname{Del}_{(r,R)}(\mathbb{R}^{d})\subsetneqq\operatorname{UD}_{r}(\mathbb{R}^{d})\subsetneqq\operatorname{QD}(\mathbb{R}^{d})\subsetneqq\mathcal{M}(\mathbb{R}^{d});$
  2. (2)
    the space $\operatorname{QD}(\mathbb{R}^{d})$ is complete and metrisable;
  3. (3)
    the space $\operatorname{UD}_{r}(\mathbb{R}^{d})$ is compact.
As a corollary of (1)–(3), $\operatorname{Del}_{(r,R)}(\mathbb{R}^{d})$ is a compact metrisable space.
The space $\operatorname{Del}_{(r,R)}(\mathbb{R}^{d})$ carries an action of $\mathbb{R}^{d}$ by translation. That is, given $\Lambda\in\operatorname{Del}_{(r,R)}(\mathbb{R}^{d})$, viewed as a discrete point set in $\mathbb{R}^{d}$; and given $a\in\mathbb{R}^{d}$, then $\Lambda+a$ is the discrete point set consisting of $x+a$ for all $x\in\Lambda$. This coincides with the $\mathbb{R}^{d}$-action on the space of measures.
From now on we shall fix a Delone set $\Lambda$. Let $\Omega_{\Lambda}$ be the closure of the orbit of $\Lambda$ under the $\mathbb{R}^{d}$-action. Then $\Omega_{\Lambda}$ is a disjoint union of orbits, hence closed under the $\mathbb{R}^{d}$-action. This allows us to restrict the topological dynamical system $\operatorname{Del}_{(r,R)}(\mathbb{R}^{d})\curvearrowleft\mathbb{R}^{d}$ to the smaller space $\Omega_{\Lambda}$ and construct the action groupoid $\Omega_{\Lambda}\rtimes\mathbb{R}^{d}$. The elements are of the form $(\omega,a)\in\Omega_{\Lambda}\times\mathbb{R}^{d}$ and the structure maps are given by
$r(\omega,a)$ $=\omega,$ $\quad s(\omega,a)$ $=\omega-a,$
$(\omega,a)\cdot(\omega-a,b)$ $=(\omega,a+b),$ $\quad(\omega,a)^{-1}$ $=(\omega-a,-a).$
[…]
Definition and Lemma 3.2 ([Bourne-Mesland:Aperiodic]*Proposition 2.14).
The action groupoid $\Omega_{\Lambda}\rtimes\mathbb{R}^{d}$ admits the following abstract transversal
$\Omega_{0}\mathrel{\vcentcolon=}\{\omega\in\Omega_{\Lambda}\mid 0\in\omega\}.$
The groupoid $\mathcal{G}_{\Lambda}$ associated to a Delone set $\Lambda$ is defined as the restriction of $\Omega_{\Lambda}\rtimes\mathbb{R}^{d}$ to the transversal $\Omega_{0}$, i.e.
$\mathcal{G}_{\Lambda}\mathrel{\vcentcolon=}\big{(}\Omega_{\Lambda}\rtimes\mathbb{R}^{d}\big{)}^{\Omega_{0}}_{\Omega_{0}}\rightrightarrows\Omega_{0}.$
Then $\mathcal{G}_{\Lambda}$ is an étale groupoid, which is Morita equivalent to $\Omega_{\Lambda}\rtimes\mathbb{R}^{d}$.
3.1.1. The regular representation
The reduced C*-algebra of an étale groupoid $\mathcal{G}$ is the completion of the groupoid convolution algebra $\mathrm{C}_{\mathrm{c}}(\mathcal{G})$ under the regular representation (cf. [Khoshkam-Skandalis:Regular_representation]), which we recall here. The real groupoid convolution algebra $\mathrm{C}_{\mathrm{c}}(\mathcal{G})$ consists of real-valued, compactly supported continuous functions $f\colon\mathcal{G}\to\mathbb{R}$, equipped with the convolution product and involution
$f*g(\gamma)\mathrel{\vcentcolon=}\sum_{\gamma_{1}\gamma_{2}=\gamma}f(\gamma_{1})g(\gamma_{2}),\quad f^{*}(\gamma)\mathrel{\vcentcolon=}f(\gamma^{-1}).$
The space $\mathrm{C}_{\mathrm{c}}(\mathcal{G})$ is a pre-Hilbert module over $\mathrm{C}_{0}(\mathcal{G}^{0})$, with right $\mathrm{C}_{0}(\mathcal{G}^{0})$-module structure and inner product (cf. [Bourne-Mesland:Aperiodic]*Section 1.3)
$(f\cdot\varphi)(\gamma)$ :$\mathrel{\vcentcolon=}f(\gamma)\varphi(s(\gamma)),$
$\left\langle f,g\right\rangle(x)$ :$\mathrel{\vcentcolon=}(f^{*}*g)\big{|}_{\mathcal{G}^{0}}(x)=\sum_{\gamma\in r^{-1}(x)}f(\gamma)g(\gamma),$
for $\gamma\in\mathcal{G}$, $x\in\mathcal{G}^{0}$, $\varphi\in\mathrm{C}_{0}(\mathcal{G}^{0})$ and $f,g\in\mathrm{C}_{\mathrm{c}}(\mathcal{G})$. Write $L^{2}(\mathcal{G})$ for the Hilbert C*-module completion of $\mathrm{C}_{\mathrm{c}}(\mathcal{G})$. Then the left multiplication
$\pi(f)\colon\mathrm{C}_{\mathrm{c}}(\mathcal{G})\to\mathrm{C}_{\mathrm{c}}(\mathcal{G}),\quad g\mapsto f*g$
extends to a *-representation
$\pi\colon\mathrm{C}_{\mathrm{c}}(\mathcal{G})\to\mathbb{B}_{\mathrm{C}_{0}(\mathcal{G}^{0})}(L^{2}(\mathcal{G})),\quad f\mapsto\pi(f),$
representing $\mathrm{C}_{\mathrm{c}}(\mathcal{G})$ by adjointable operators on Hilbert $\mathrm{C}_{0}(\mathcal{G}^{0})$-module $L^{2}(\mathcal{G})$.
Definition 3.3.
The reduced groupoid C*-algebra $\mathrm{C}^{*}(\mathcal{G})$ of an étale groupoid $\mathcal{G}$ is the completion of $\mathrm{C}_{\mathrm{c}}(\mathcal{G})$ in the norm $f\mapsto\left\lVert\pi(f)\right\rVert$.
Let $x\in\mathcal{G}^{0}$. Then the evaluation *-homomorphism
$\operatorname{ev}_{x}\colon\mathrm{C}_{0}(\mathcal{G}^{0})\to\mathbb{R},\quad\operatorname{ev}_{x}(\varphi)\mathrel{\vcentcolon=}\varphi(x)$
induces a *-homomorphism
$(\operatorname{ev}_{x})_{*}\colon\mathbb{B}_{\mathrm{C}_{0}(\mathcal{G}^{0})}(L^{2}(\mathcal{G}))\to\mathbb{B}(L^{2}(\mathcal{G})\otimes_{\operatorname{ev}_{x}}\mathbb{R}),\quad T\mapsto T\otimes\operatorname{id}.$
Denote by $\mathcal{G}_{x}$ the source fibre of $\mathcal{G}$ at $x$. The Hilbert space
$\mathcal{H}_{x}\mathrel{\vcentcolon=}L^{2}(\mathcal{G})\otimes_{\operatorname{ev}_{x}}\mathbb{R}$
is isomorphic to $\ell^{2}(\mathcal{G}_{x})$, sending $f\otimes t$ to the restriction of $t\cdot f$ to $\mathcal{G}_{x}$. Thus $\pi_{x}\mathrel{\vcentcolon=}(\operatorname{ev}_{x})_{*}\circ\pi$ gives a representation of $\mathrm{C}^{*}(\mathcal{G})$ on $\mathcal{H}_{x}$. We call it the localised representation (of the regular representation $\pi$) at $x\in\mathcal{G}^{0}$; and call $\mathcal{H}_{x}$ the localised Hilbert space at $x\in\mathcal{G}^{0}$.
Now we let $\mathcal{G}$ be the groupoid of Delone sets $\mathcal{G}_{\Lambda}$. By definition, its regular representation is defined on the Hilbert $\mathrm{C}(\Omega_{0})$-module $L^{2}(\mathcal{G}_{\Lambda})_{\mathrm{C}(\Omega_{0})}$. This is a Hilbert module over a unital, commutative C*-algebra, and hence equivalent to a continuous field of Hilbert spaces over $\Omega_{0}$. The source fibre of $\mathcal{G}_{\Lambda}$ at $\omega\in\Omega_{0}$ is given by
[…]
which is in bijection with the Delone set $\omega$ via $(\omega-x,-x)\mapsto x$. Thus the localised Hilbert space $\mathcal{H}_{\omega}$ is unitarily isomorphic to $\ell^{2}(\omega)$ via (cf. [Bourne-Mesland:Aperiodic]*Section 4.1):
(3.4) $\rho_{\omega}\colon\mathcal{H}_{\omega}\xrightarrow{\sim}\ell^{2}(s^{-1}(\omega)),$ $\qquad\rho_{\omega}(f\otimes t)(x)\mathrel{\vcentcolon=}tf(\omega-x,-x).$
$\rho_{\omega}^{-1}\colon\ell^{2}(s^{-1}(\omega))\xrightarrow{\sim}\mathcal{H}_{\omega},$ $\qquad\rho_{\omega}^{-1}(\left\lvert x\right\rangle)\mathrel{\vcentcolon=}\left\lvert\omega-x,-x\right\rangle,$
where $\left\lvert\omega-x,-x\right\rangle$ is the equivalence class in $\mathcal{H}_{\omega}$ of any continuous function $f$ on $\mathcal{G}$, such that $\operatorname{supp}f\cap s^{-1}(\omega)=\{(\omega-x,-x)\}$ and $f(\omega-x,-x)=1$.
Lemma 3.5 ([Bourne-Mesland:Aperiodic]*Section 4.1).
The localised representation $\pi_{\omega}\colon\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\to\mathbb{B}(\ell^{2}(\omega))$ is given by the formula
(3.6) $(\pi_{\omega}(f)\psi)(x)\mathrel{\vcentcolon=}\sum_{y\in\omega}f(\omega-x,y-x)\psi(y)$
for $f\in\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$, $\psi\in\ell^{2}(\omega)$ and $x\in\omega$. Therefore,
$\langle x\,|\pi_{\omega}(f)|\,\mathopen{}y\rangle=f(\omega-x,y-x),\quad x,y\in\omega.$
When we consider systems with a finite number of internal degrees of freedom inside every lattice site, e.g. spins of electrons, then we must replace the observable C*-algebra by its matrix algebra $\mathbb{M}_{N}\mathrm{C}^{*}(\mathcal{G}_{\Lambda})=\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\otimes\mathbb{M}_{N}(\mathbb{R})$. We extend the regular representation of $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$ to a representation
$\pi_{\omega}^{N}\mathrel{\vcentcolon=}\pi_{\omega}\otimes e_{N},\quad\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\otimes\mathbb{M}_{N}(\mathbb{R})\to\mathbb{B}(\ell^{2}(\omega,\mathcal{K}))$
where $\mathcal{K}$ is a separable Hilbert space, and $e_{N}\colon\mathbb{M}_{N}(\mathbb{R})\hookrightarrow\mathbb{K}(\mathcal{K})$ is any rank-$N$ corner embedding. We may describe the representation $\pi_{\omega}^{N}$ by an infinite matrix indexed by $x,y\in\omega$, with matrix elements
(3.7) $\pi^{N}_{\omega}(f\otimes S)_{x,y}\mathrel{\vcentcolon=}\left\langle x\,\middle|\pi_{\omega}(f)\middle|\,\mathopen{}y\right\rangle\cdot S\in\mathbb{M}_{N}(\mathbb{R}).$
for all $f\in\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$ and $S\in\mathbb{M}_{N}(\mathbb{R})$.
3.1.2. Position spectral triple over the groupoid C*-algebra
In order to construct a spectral triple over $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$ or $\mathbb{M}_{N}\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$, one way is to first construct an unbounded Kasparov $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$-$\mathrm{C}(\Omega_{0})$–module, then take its (unbounded) Kasparov product with a *-homomorphism $\mathrm{C}(\Omega_{0})\to\mathbb{R}$. Such an unbounded Kasparov module was given in [Bourne-Mesland:Aperiodic]*Section 2.3.1 using the “position” operators on the Hilbert $\mathrm{C}(\Omega_{0})$-module $L^{2}(\mathcal{G}_{\Lambda})$:
(3.8) $\prescript{}{d}{\lambda}_{\Omega_{0}}\mathrel{\vcentcolon=}\left(\mathrm{C}_{\mathrm{c}}(\mathcal{G}_{\Lambda})\otimes\mathrm{C}\ell_{0,d},\quad L^{2}(\mathcal{G}_{\Lambda})_{\mathrm{C}(\Omega_{0})}\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d},\quad\sum_{i=1}^{d}X_{j}\otimes\gamma^{j}\right),$
where
$c_{j}\colon\mathcal{G}_{\Lambda}\to\mathbb{R},\quad c_{j}(\omega,x)\mathrel{\vcentcolon=}x_{j};\qquad X_{j}f(\omega,x)\mathrel{\vcentcolon=}c_{j}(\omega,x)f(\omega,x).$
The map $c_{j}\colon\mathcal{G}_{\Lambda}\to\mathbb{R}$ is an exact groupoid cocycle in the sense of [Mesland:Groupoid_cocycles]*Definition 4.1.2. It follows from the general construction in [Mesland:Groupoid_cocycles]*Theorem 3.2.2 that $\prescript{}{d}{\lambda}_{\Omega_{0}}$ is an unbounded Kasparov module. Following [Bourne-Mesland:Aperiodic], we call $\prescript{}{d}{\lambda}_{\Omega_{0}}$ the bulk cycle.
[…]
Theorem 3.9.
Let $\omega\in\Omega_{0}$ and $\mathcal{K}$ be a Hilbert space. The following data
(3.10) $\xi_{\omega,N}^{\mathrm{Gpd}}\mathrel{\vcentcolon=}\left(\mathbb{M}_{N}(\mathrm{C}_{\mathrm{c}}(\mathcal{G}_{\Lambda}))\otimes\mathrm{C}\ell_{0,d},\quad\ell^{2}(\omega,\mathcal{K})\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d},\quad\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}\right)$
is a locally compact position spectral triple. It represents a class in
$\operatorname{KK}(\mathbb{M}_{N}\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\otimes\mathrm{C}\ell_{0,d},\mathbb{R})\simeq\operatorname{KK}(\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\otimes\mathrm{C}\ell_{0,d},\mathbb{R}),$
which is the Kasparov product of the following unbounded Kasparov modules and *-homomorphisms:
  1. (1)
    the bulk cycle $\prescript{}{d}{\lambda}_{\Omega_{0}}$, which gives a class in $\operatorname{KK}(\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\otimes\mathrm{C}\ell_{0,d},\mathrm{C}(\Omega_{0}))$;
  2. (2)
    the evaluation *-homomorphism $\operatorname{ev}_{\omega}\colon\mathrm{C}(\Omega_{0})\to\mathbb{R}$, which gives a class in $\operatorname{KK}(\mathrm{C}(\Omega_{0}),\mathbb{R})$;
  3. (3)
    exterior product with the imprimitivity bimodule $(\mathbb{M}_{N}(\mathbb{R}),\mathcal{K},0)$, which represnts the identity element of $\operatorname{KK}(\mathbb{R},\mathbb{R})$.
Proof.
The composition of (1)–(3) is easy to compute because either (2) or (3) can be described by a Hilbert module with zero operator. The tensor product of these Hilbert modules
$\left(L^{2}(\mathcal{G}_{\Lambda})\otimes_{\operatorname{ev}_{\omega}}\mathbb{R}\otimes\mathcal{K}\right)\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d}$
is isomorphic to $\ell^{2}(\omega,\mathcal{K})\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d}$ via the localised representation at $\omega$. Thus it suffices to check that $c_{j}$ is mapped to ${\scriptstyle\mathsf{X}}_{j}$ under the composition and that (3.10) is a spectral triple. By (3.4), the operator $X_{j}\otimes\operatorname{id}_{\mathcal{K}}\otimes\operatorname{id}_{\bigwedge\nolimits^{*}\mathbb{R}^{d}}$ acts on $\ell^{2}(\omega,\mathcal{K})\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d}$ sends $\left\lvert x\right\rangle\otimes v\otimes w$ to $x_{j}\left\lvert x\right\rangle\otimes v\otimes w$, where $x\in\omega$ is considered an element in $\mathbb{R}^{d}$. It is a spectral triple by [Bourne-Mesland:Aperiodic]*Proposition 4.1. Thus (3.10) is a position spectral triple. It is locally compact in the sense of Definition 2.16 because the matrix elements of $\pi^{N}_{\omega}$ belong to $\mathbb{M}_{N}(\mathbb{R})$, hence are compact operators on $\mathcal{K}$. ∎
3.2. Roe C*-algebra of a Delone set
The coarse-geometric approach uses the Roe C*-algebra $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$ of the Delone set $\Lambda$. It consists of operators on $\ell^{2}(\Lambda,\mathcal{K})$ which are locally compact and can be approximated by controlled operators. In this setting, we treat a Delone set $\Lambda\subseteq\mathbb{R}^{d}$ as discrete metric space, equipped with the subspace topology from $\mathbb{R}^{d}$.
Definition 3.11.
Let $X$ be a proper metric space. An ample $X$-module or ample module over $X$ is given by a pair $(X,\mathcal{H}_{X})$, where $X$ is a proper metric space, and $\mathcal{H}_{X}$ is a Hilbert space that carries a non-degenerate representation $\varrho\colon\mathrm{C}_{0}(X)\to\mathbb{B}(\mathcal{H}_{X})$, which is ample in the sense that
$\varrho(f)\in\mathbb{K}(\mathcal{H}_{X})$ iff $f=0$.
Example 3.12 ([Willett-Yu:Higher_index_theory]*Example 4.1.2–4.1.4).
Let $X$ be a Riemannian manifold with volume form $\mu$. Assume that every connected component of $X$ has dimension greater equal than 1, e.g. $X=\mathbb{R}^{d}$ for $d\geq 1$ with the Lebesgue measure. Then pointwise multiplication gives an ample $X$-module $\mathcal{H}_{X}\mathrel{\vcentcolon=}L^{2}(X,\mu)$.
If $X$ is a discrete metric space with the counting measure, then the representation $\mathrm{C}_{0}(X)\to\mathbb{B}(\ell^{2}(X))$ by multiplication is no longer ample. In such cases, we replace $\ell^{2}(X)$ by by $\mathcal{H}_{X}\mathrel{\vcentcolon=}\ell^{2}(X,\mathcal{K})=\ell^{2}(X)\otimes\mathcal{K}$, where $\mathcal{K}$ is any separable Hilbert space, and the representation maps $f\in\mathrm{C}_{0}(X)$ to the tensor product of pointwise multiplication with $f$ with the identity operator on $\mathcal{K}$. Then $\mathcal{H}_{X}$ is an ample $X$-module.
We call $\mathcal{H}_{X}$ in the above cases, the standard ample $X$-module.
Definition 3.13 ([Willett-Yu:Higher_index_theory]*Definition 4.1.7 and 5.1.1).
Let $(X,\mathcal{H}_{X})$ and $(Y,\mathcal{H}_{Y})$ be ample modules, carrying ample representations $\varrho_{X}\colon\mathrm{C}_{0}(X)\to\mathbb{B}(\mathcal{H}_{X})$ and $\varrho_{Y}\colon\mathrm{C}_{0}(Y)\to\mathbb{B}(\mathcal{H}_{Y})$.
  • The support of an operator $T\in\mathbb{B}(\mathcal{H}_{X},\mathcal{H}_{Y})$, denoted by $\operatorname{supp}(T)$, is the collection of all points $(y,x)\in Y\times X$, such that $\varrho_{Y}(\chi_{V})T\varrho_{X}(\chi_{U})\neq 0$ holds for all open neighbourhood $U$ of $x$ and $V$ of $y$.
  • A subset $\mathcal{E}\subseteq X\times X$ is called controlled, if
    $\sup\left\{\mathrm{d}_{X}(x_{1},x_{2})\;\middle|\;(x_{1},x_{2})\in\mathcal{E}\right\}<+\infty.$
  • An operator $S\in\mathbb{B}(\mathcal{H}_{X})$ is called locally compact, if for any $f\in\mathrm{C}_{\mathrm{c}}(X)$, $T\varrho(f)$ and $\varrho(f)T$ are compact operators.
  • An operator $S\in\mathbb{B}(\mathcal{H}_{X})$ is called controlled, if $\operatorname{supp}(S)$ is controlled.
Definition 3.14 ([Willett-Yu:Higher_index_theory]*Definition 5.1.4).
Let $(X,\mathcal{H}_{X})$ be an ample module. The Roe C*-algebra $\mathrm{C}^{*}_{\mathrm{Roe}}(X,\mathcal{H}_{X})$ is the C*-algebra generated by all locally compact, controlled operators on $\mathcal{H}_{X}$.
Example 3.15 ([Ewert-Meyer:Coarse_geometry]*Example 1).
Let $X$ be a discrete metric space and $\mathcal{H}_{X}=\ell^{2}(X,\mathcal{K})$ be the standard ample $X$-module. Describe an operator $T\in\mathbb{B}(\mathcal{H}_{X})$ by an infinite matrix $(T_{x,y})_{x,y\in X}$ where
$T_{x,y}\mathrel{\vcentcolon=}\left\langle x\,\middle|T\middle|\,\mathopen{}y\right\rangle\in\mathbb{B}(\mathcal{K}).$
Then the support of $T$ is given by
$\operatorname{supp}(T)=\left\{(x,y)\in X\times X\;\middle|\;T_{x,y}\neq 0\right\}.$
The operator $T$ is locally compact, iff $T_{x,y}\in\mathbb{K}(\mathcal{K})$ for all $x,y\in X$; controlled, iff there exists $R>0$, such that $T_{x,y}=0$ whenever $\mathrm{d}(x,y)>R$.
A *-homomorphism between Roe C*-algebras can be induced by a certain (not necessarily continuous) map between their underlying spaces. Such maps are called coarse. Their induced *-homomorphism are constructed using covering isometries.
Definition 3.16 ([Willett-Yu:Higher_index_theory]*Definition 5.1.10).
Let $(X,\mathrm{d}_{X})$ and $(Y,\mathrm{d}_{Y})$ be proper metric spaces. A map $f\colon X\to Y$ is called coarse, if the following holds:
  1. (1)
    The expansion function of $f$, given by
    $\omega_{f}\colon[0,+\infty)\to[0,+\infty),\quad\omega_{f}(r)\mathrel{\vcentcolon=}\left\{\mathrm{d}_{Y}(f(x_{1}),f(x_{2}))\;\middle|\;\mathrm{d}_{X}(x_{1},x_{2})\leq r\right\},$
    satisfies $\omega_{f}(r)<+\infty$ for all $r\geq 0$.
  2. (2)
    The map $f$ is proper, i.e. $f^{-1}(K)\subseteq X$ is pre-compact for any compact set $K\subseteq Y$.
Two coarse maps $f,g\colon X\rightrightarrows Y$ are close, if there exists $c\geq 0$ such that for all $x\in X$, $d_{Y}(f(x),g(x))\leq c$.
Definition and Lemma 3.17 ([Willett-Yu:Higher_index_theory]*Definition 5.1.11, Lemma 5.1.12).
Let $(X,\mathcal{H}_{X})$ and $(Y,\mathcal{H}_{Y})$ be ample modules and let $f\colon X\to Y$ be a coarse map. A covering isometry for $f$ is an isometry $V\colon\mathcal{H}_{X}\to\mathcal{H}_{Y}$ such that there exists $t\geq 0$ such that $\mathrm{d}(y,f(x))<t$ whenever $(y,x)\in\operatorname{supp}(V)$.
The *-homomorphism
$\operatorname{Ad}_{V}\colon\mathbb{B}(\mathcal{H}_{X})\to\mathbb{B}(\mathcal{H}_{Y}),\qquad T\mapsto VTV^{*}$
restricts to a *-homomorphism $\mathrm{C}^{*}_{\mathrm{Roe}}(X,\mathcal{H}_{X})\to\mathrm{C}^{*}_{\mathrm{Roe}}(Y,\mathcal{H}_{Y})$. Its induced maps
$f_{*}\colon\operatorname{KK}(\mathrm{C}\ell_{k,0},\mathrm{C}^{*}_{\mathrm{Roe}}(X,\mathcal{H}_{X}))\to\operatorname{KK}(\mathrm{C}\ell_{k,0},\mathrm{C}^{*}_{\mathrm{Roe}}(Y,\mathcal{H}_{Y}))$
depends only on $f$ and not on the choice of $V$.
The coarse category consists of proper metric spaces as objects, and equivalence classes of coarse maps as arrows, where two coarse maps are equivalent iff they are close. A coarse equivalence is an isomorphism in the coarse category. If $X\subseteq Y$ is a metric subspace that is coarsely equivalent to $Y$, then we also say that $X$ is coarsely dense in $Y$. In particular, any Delone set $\Lambda$ of $\mathbb{R}^{d}$ is coarsely dense in $\mathbb{R}^{d}$.
Proposition 3.18 ([Willett-Yu:Higher_index_theory]*Proposition 4.3.5).
Let $(X,\mathcal{H}_{X})$ and $(Y,\mathcal{H}_{Y})$ be ample modules and $f\colon X\to Y$ be a coarse equivalence, then there exists a covering isometry $V\colon\mathcal{H}_{X}\to\mathcal{H}_{Y}$ that is a unitary equivalence.
As a special case, we have the following:
Proposition 3.19.
Let $Y$ be a discrete metric space, and $X\subseteq Y$ be a coarsely dense subset. Let $\mathcal{H}_{X}\mathrel{\vcentcolon=}\ell^{2}(X,\mathcal{K})$ and $\mathcal{H}_{Y}\mathrel{\vcentcolon=}\ell^{2}(Y,\mathcal{K})$ be their standard ample modules. Then the operator $V\colon\mathcal{H}_{X}\to\mathcal{H}_{Y}$, $V\left\lvert x\right\rangle\mathrel{\vcentcolon=}\left\lvert x\right\rangle$ is a covering isometry for the isometric embedding $\iota\colon X\hookrightarrow Y$ and induce isomorphisms
$\iota_{*}\colon\operatorname{KK}(\mathrm{C}\ell_{k,0},\mathrm{C}^{*}_{\mathrm{Roe}}(X,\mathcal{H}_{X}))\to\operatorname{KK}(\mathrm{C}\ell_{k,0},\mathrm{C}^{*}_{\mathrm{Roe}}(Y,\mathcal{H}_{Y})).$
Proof.
The support of $V$ is
$\operatorname{supp}(V)=\left\{(y,x)\in Y\times X\;\middle|\;\left\langle y\,\middle|V\middle|\,\mathopen{}x\right\rangle\neq 0\right\}=\left\{(\iota(x),x)\;\middle|\;x\in X\right\}.$
So $\mathrm{d}(y,\iota(x))=0$ for all $(y,x)\in\operatorname{supp}(V)$. Therefore $V$ is an covering isometry for $\iota$, and $\operatorname{Ad}_{V}$ induces a *-homomorphism $\iota_{*}\colon\mathrm{C}^{*}_{\mathrm{Roe}}(X,\mathcal{H}_{X})\to\mathrm{C}^{*}_{\mathrm{Roe}}(Y,\mathcal{H}_{Y})$. Since $\iota$ is a coarse equivalence, there exists a covering isometry $V^{\prime}\colon\mathcal{H}_{X}\to\mathcal{H}_{Y}$ that is a unitary equivalence. Hence, $\operatorname{Ad}_{V^{\prime}}$ induces an isomorphism
$\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(X,\mathcal{H}_{X}))\to\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(Y,\mathcal{H}_{Y})).$
By 3.17, the induced maps of $\operatorname{Ad}_{V}$ coincides with that of $\operatorname{Ad}_{V^{\prime}}$, hence an isomorphism as well. ∎
Remark 3.20.
By Proposition 3.19, the Roe C*-algebra of a Delone set $\Lambda\subseteq\mathbb{R}^{d}$ is unique up to isomorphism, independent of its defining ample module. In particular, it is isomorphic to the Roe C*-algebra of $\mathbb{R}^{d}$. However, we must be careful as the inclusion $\iota\colon\Lambda\hookrightarrow\mathbb{R}^{d}$ does not induce an isomorphism (via a covering isometry) between the corresponding Roe C*-algebras $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda,\mathcal{H}_{\Lambda})$ and $\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{R}^{d},\mathcal{H}_{\mathbb{R}^{d}})$ defined using their standard ample modules. An isomorphism between the Roe C*-algebras of coarsely equivalent spaces $X$ and $Y$ was constructed in [Ewert-Meyer:Coarse_geometry]*Theorem 3. However, these spaces carry different ample modules, involving an inverse of the coarsely dense embedding $\iota\colon X\to X\coprod Y$ in the coarse category.
As opposed to the groupoid C*-algebra $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$, the Roe C*-algebra $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$ has simple K-theory groups. It follows from coarse Mayer–Vietoris sequence (cf. [Higson-Roe-Yu:Coarse_Mayer-Vietoris]) that the K-theory of $\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{R}^{d})$ coincides with the K-theory of a point by a degree shift of $-d$. This holds for any Delone set $\Lambda\subseteq\mathbb{R}^{d}$ as well by Proposition 3.19. Therefore, K-theory of the real or complex Roe C*-algebras of a Delone set $\Lambda\subseteq\mathbb{R}^{d}$ is given by
(3.21) $\operatorname{K}_{i}(\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)_{\mathbb{C}})$$\simeq\begin{cases}\mathbb{Z}{\hphantom{/2}}&\text{if }i-d\equiv 0{\hphantom{,5,6,7}}\mkern-10.0mu\mod 2;\\ 0&\text{if }i-d\equiv 1{\hphantom{,5,6,7}}\mkern-10.0mu\mod 2.\end{cases}$
$\operatorname{K}_{i}(\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)_{\mathbb{R}})$$\simeq\begin{cases}\mathbb{Z}&\text{if }i-d\equiv 0,4{\hphantom{,6,7}}\mkern-10.0mu\mod 8;\\ \mathbb{Z}/2&\text{if }i-d\equiv 1,2{\hphantom{,6,7}}\mkern-10.0mu\mod 8;\\ 0&\text{if }i-d\equiv 3,5,6,7\mkern-10.0mu\mod 8.\end{cases}$
While using the Roe C*-algebra as the observable C*-algebra, there is no need to pass to its matrix algebras, as opposed to the case of groupoids. It was proven in [Ewert-Meyer:Coarse_geometry]*Corollary 1 that an operator $T$ on $\mathcal{H}_{\Lambda}^{\oplus N}$ is locally compact and controlled iff its matrix elements in $\mathbb{B}(\mathcal{H}_{\Lambda})$ are locally compact and controlled, when $T$ is viewed as an $N\times N$-matrix with entries in $\mathbb{B}(\mathcal{H}_{\Lambda})$. Therefore
$\mathbb{M}_{N}\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda,\mathcal{H}_{\Lambda})\simeq\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda,\mathcal{H}_{\Lambda}^{\oplus N})$
for all $N$, where $\mathcal{H}_{\Lambda}^{\oplus N}$ carries the direct sum of the standard ample representation $\mathrm{C}_{0}(\Lambda)\to\mathbb{B}(\mathcal{H}_{\Lambda})$. We may further identify $\mathcal{H}_{\Lambda}^{\oplus}\simeq\ell^{2}(\Lambda,\mathcal{K})\otimes\mathbb{R}^{N}\simeq\ell^{2}(\Lambda,\mathcal{K}\otimes\mathbb{R}^{N})\simeq\ell^{2}(\Lambda,\mathcal{K})$ using any unitary isomorphism $\mathcal{K}\simeq\mathcal{K}\otimes\mathbb{R}^{N}$. Such an isomorphism preserves controlled or locally compact operators using the characterisation in Example 3.15.
3.2.1. Position spectral triple over the Roe C*-algebra
Fix a closed subset $X\subseteq\mathbb{R}^{d}$ equipped with the subspace metric, e.g. a Delone set in $\mathbb{R}^{d}$. A relation between the Roe C*-algebra on $X$ and the “position” operators has been considered in [Ewert-Meyer:Coarse_geometry]*Section 2.2 as follows. For $t=(t_{1},\dots,t_{d})\in\mathbb{R}^{d}$, let $\mathrm{e}_{t}$ be the bounded continuous function on $X\subseteq\mathbb{R}^{d}$ given by
$\mathrm{e}_{t}(x)\mathrel{\vcentcolon=}\mathrm{e}^{\mathrm{i}t\cdot x}=\mathrm{e}^{\mathrm{i}\sum_{j=1}^{d}t_{j}x_{j}},\qquad x=(x_{1},\dots,x_{d})\in X\subseteq\mathbb{R}^{d}.$
Then the map $\mathbb{R}^{d}\to\mathrm{C}_{\mathrm{b}}(X)$ given by $t\mapsto\mathrm{e}_{t}$ is continuous in the strict topology of $\mathrm{C}_{\mathrm{b}}(X)$.
Let $\mathcal{H}_{X}$ be an ample $X$-module. The representation $\varrho\colon\mathrm{C}_{0}(X)\to\mathbb{B}(\mathcal{H}_{X})$ extends to a strictly continuous, *-representation $\overline{\varrho}\colon\mathrm{C}_{\mathrm{b}}(X)\to\mathbb{B}(\mathcal{H}_{X})$. Thus the map
$\sigma\colon\mathbb{R}^{d}\to\mathrm{C}_{\mathrm{b}}(X)\to\mathbb{B}(\mathcal{H}_{X}),\qquad t\mapsto\mathrm{e}_{t}\mapsto\varrho(\mathrm{e}_{t})$
is continuous for the norm topology on $\mathbb{B}(\mathcal{H}_{X})$. If $X$ does not contain discrete components, and $\mathcal{H}_{X}$ is the standard $X$-module as in Example 3.12, then the restriction of $\sigma$ to the $j$-th coordinate component of $\mathbb{R}^{d}$ gives the flow $\mathrm{e}^{\mathrm{i}t{\scriptstyle\mathsf{X}}_{j}}$ generated by ${\scriptstyle\mathsf{X}}_{j}$.
By conjugation, $\sigma\colon\mathbb{R}^{d}\to\mathbb{B}(\mathcal{H}_{X})$ induces a group homomorphism $\operatorname{Ad}\sigma\colon\mathbb{R}^{d}\to\operatorname{Aut}(\mathbb{B}(\mathcal{H}_{X}))$. We say that an operator $T\in\mathbb{B}(\mathcal{H}_{X})$ is continuous with respect to $\operatorname{Ad}\sigma$ iff the map $t\mapsto\operatorname{Ad}\sigma_{t}(T)$ is continuous for the norm topology on $\mathbb{B}(\mathcal{H})$. The property of being a norm limit of controlled operators can be described as a continuity property for the action $\operatorname{Ad}\sigma$:
Lemma 3.22 ([Ewert-Meyer:Coarse_geometry]*Theorem 4).
An operator $T\in\mathbb{B}(\mathcal{H}_{X})$ is a norm limit of controlled operators iff it is continuous with respect to $\operatorname{Ad}\sigma$. Thus $T\in\mathrm{C}^{*}_{\mathrm{Roe}}(X,\mathcal{H}_{X})$ iff it is locally compact and continuous with respect to $\operatorname{Ad}\sigma$.
We take a slight different viewpoint, interpreting the previous lemma as follows: the Roe C*-algebra is the largest C*-algebra, over which one can have a locally compact position spectral triple:
Theorem 3.23.
Let $\Lambda\subseteq\mathbb{R}^{d}$ be a discrete subset. If
$\left(\mathcal{A}\otimes\mathrm{C}\ell_{0,d},\quad\ell^{2}(\Lambda,\mathcal{K})\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d},\quad\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}\right)$
is a locally compact position spectral triple (cf. Definition 2.16), where $\mathcal{A}$ is a dense *-subalgebra of a C*-algebra $A$ represented on $\ell^{2}(\Lambda,\mathcal{K})$ by $\varphi\colon A\to\mathbb{B}(\ell^{2}(\Lambda,\mathcal{K}))$. Then $\varphi(\mathcal{A})$ is contained in the Roe C*-algebra $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda,\mathcal{H}_{\Lambda})$ defined by the standard $\Lambda$-module $\mathcal{H}_{\Lambda}\mathrel{\vcentcolon=}\ell^{2}(\Lambda,\mathcal{K})$ as in Example 3.12.
Proof.
Identifying $A$ and $\mathcal{A}$ with their images under $\varphi$, we may assume that $\varphi\colon A\to\mathbb{B}(\ell^{2}(\Lambda,\mathcal{K}))$ is injective. Then we must show that every $a\in\mathcal{A}$ is locally compact and can be approximated by controlled operators. By definition of a locally compact spectral triple, it follows that $\left\langle x\,\middle|a\middle|\,\mathopen{}y\right\rangle\in\mathbb{K}(\mathcal{K})$ for all $x,y\in\Lambda$. This is the same condition for $a\in\mathcal{A}$ being locally compact in the sense of Definition 3.13.
Now we show that every $a\in\mathcal{A}$ is a norm limit of controlled operators. Since $[a\otimes\rho^{i},{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}]$ is bounded for all $a\in\mathcal{A}$ and $i,j\in\{1,\dots,d\}$, we have $[a,{\scriptstyle\mathsf{X}}_{j}]$ is a bounded operator for all $j\in\{1,\dots,d\}$, i.e. there exists $C>0$ such that $\left\lVert[a,{\scriptstyle\mathsf{X}}_{j}]\right\rVert\leq C$.
By the previous lemma, it suffices to show that for all $a\in\mathcal{A}$ and $j=1,\dots,d$, the map
$\mathbb{R}\to\mathbb{B}(\ell^{2}(\Lambda,\mathcal{K})),\qquad t\mapsto\mathrm{e}^{\mathrm{i}t{\scriptstyle\mathsf{X}}_{j}}a\mathrm{e}^{-\mathrm{i}t{\scriptstyle\mathsf{X}}_{j}}$
is continuous in the norm topology on $\mathbb{B}(\ell^{2}(\Lambda,\mathcal{K}))$. It follows from
$[\mathrm{e}^{it{\scriptstyle\mathsf{X}}_{j}},a]=\int_{0}^{1}\frac{\mathrm{d}}{\mathrm{d}s}\left(\mathrm{e}^{\mathrm{i}st{\scriptstyle\mathsf{X}}_{j}}a\mathrm{e}^{\mathrm{i}(1-s)t{\scriptstyle\mathsf{X}}_{j}}\right)\mathrm{d}s=\mathrm{i}t\int_{0}^{1}\mathrm{e}^{\mathrm{i}st{\scriptstyle\mathsf{X}}_{j}}[{\scriptstyle\mathsf{X}}_{j},a]\mathrm{e}^{\mathrm{i}(1-s)t{\scriptstyle\mathsf{X}}_{j}}\mathrm{d}s$
that
$\left\lVert[\mathrm{e}^{\mathrm{i}t{\scriptstyle\mathsf{X}}_{j}},a]\right\rVert\leq t\int_{0}^{1}\left\lVert[{\scriptstyle\mathsf{X}}_{j},a]\right\rVert\mathrm{d}s\leq Ct.$
Therefore,
$\left\lVert\mathrm{e}^{\mathrm{i}s{\scriptstyle\mathsf{X}}_{j}}a\mathrm{e}^{-\mathrm{i}s{\scriptstyle\mathsf{X}}_{j}}-\mathrm{e}^{\mathrm{i}t{\scriptstyle\mathsf{X}}_{j}}a\mathrm{e}^{-\mathrm{i}t{\scriptstyle\mathsf{X}}_{j}}\right\rVert$ $=\left\lVert\mathrm{e}^{\mathrm{i}t{\scriptstyle\mathsf{X}}_{j}}\left(\mathrm{e}^{\mathrm{i}(s-t){\scriptstyle\mathsf{X}}_{j}}a\mathrm{e}^{-\mathrm{i}(s-t){\scriptstyle\mathsf{X}}_{j}}-a\right)\mathrm{e}^{-\mathrm{i}t{\scriptstyle\mathsf{X}}_{j}}\right\rVert$
$=\left\lVert\mathrm{e}^{\mathrm{i}t{\scriptstyle\mathsf{X}}_{j}}[\mathrm{e}^{\mathrm{i}(s-t){\scriptstyle\mathsf{X}}_{j}},a]\mathrm{e}^{-\mathrm{i}(s-t){\scriptstyle\mathsf{X}}_{j}}\mathrm{e}^{-\mathrm{i}t{\scriptstyle\mathsf{X}}_{j}}\right\rVert$
$\leq\left\lVert[\mathrm{e}^{\mathrm{i}(s-t){\scriptstyle\mathsf{X}}_{j}},a]\right\rVert$
$\leq C(s-t).$
So the map $t\mapsto\operatorname{Ad}{\sigma_{t}}(a)$ is continuous for all $a\in\mathcal{A}$, and thus $a\in\mathcal{A}$ is a norm limit of controlled operators. ∎
Thus if the *-algebra $\mathcal{A}$ in the position spectral triple is dense in $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$, then it defines a spectral triple over $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$. This amounts to choosing a suitable dense *-subalgebra inside $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$. The spectral triple below has been implicitly used by Ewert and Meyer in the proof of [Ewert-Meyer:Coarse_geometry]*Theorem 7.
Definition and Lemma 3.24.
Let $X\subseteq\mathbb{R}^{d}$ be a discrete countable set. Let $\mathcal{C}_{\mathrm{Roe}}(X)$ be the collection of operators $T\in\mathbb{B}(\ell^{2}(X,\mathcal{K}))$ satisfying:
  • $T$ is controlled and locally compact;
  • $\sup_{y\in X}\sum_{x\in X}\left\lVert T_{x,y}\right\rVert<+\infty$ and $\sup_{x\in X}\sum_{y\in X}\left\lVert T_{x,y}\right\rVert<+\infty$.
Then
(3.25) $\xi^{\mathrm{Roe}}_{X}\mathrel{\vcentcolon=}\left(\mathcal{C}_{\mathrm{Roe}}(X)\otimes\mathrm{C}\ell_{0,d},\quad\ell^{2}(X,\mathcal{K})\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d},\quad\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}\right)$
is a locally compact position spectral triple over $\mathrm{C}^{*}_{\mathrm{Roe}}(X)$. If $X$ is a Delone set, then $\xi^{\mathrm{Roe}}_{X}$ is $d$-summable.
Proof.
Let $T\in\mathcal{C}_{\mathrm{Roe}}(X)$, $x,y\in X$. Then
$\langle x\,|[T,{\scriptstyle\mathsf{X}}_{j}]|\,\mathopen{}y\rangle=\langle x\,|T{\scriptstyle\mathsf{X}}_{j}|\,\mathopen{}y\rangle-\langle x\,|{\scriptstyle\mathsf{X}}_{j}T|\,\mathopen{}y\rangle=(y_{j}-x_{j})T_{x,y}.$
Since $T$ is controlled, there exists $R>0$ such that $\langle x\,|T|\,\mathopen{}y\rangle=0$ whenever $\mathrm{d}(x,y)>R$. Thus
$\left\lVert[T,{\scriptstyle\mathsf{X}}_{j}]\right\rVert\leq\sup_{y\in X}\sum_{x\in X}R\cdot\left\lVert T_{x,y}\right\rVert<+\infty.$
This shows that $T$ maps the domain of ${\scriptstyle\mathsf{X}}_{j}$ into its domain and has bounded commutator with ${\scriptstyle\mathsf{X}}_{j}$ for all $j=1,\dots,d$.
Therefore, for any $S\in\mathrm{C}\ell_{0,d}$, $T\otimes S$ maps the domain of $\sum_{j=1}^{n}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}$ into the domain of $\sum_{j=1}^{n}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}$ and
$\left[T\otimes S,\quad\sum_{j=1}^{n}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}\right]=\sum_{j=1}^{n}\left[T,{\scriptstyle\mathsf{X}}_{j}\right]\otimes\left[S,\gamma^{j}\right]$
is bounded. The operator
$\left(1+\left(\sum_{j=1}^{n}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}\right)^{2}\right)^{-1}=\left(1+\sum_{j=1}^{n}{\scriptstyle\mathsf{X}}_{j}^{2}\otimes 1\right)^{-1}$
is compact. Therefore, $\xi^{\mathrm{Roe}}_{X}$ is a spectral triple, which is a locally compact position spectral triple because $T$ is locally compact.
If $X$ is a Delone set, then it is uniformly discrete, which implies that
$\operatorname{tr}\left(1+\left(\sum_{j=1}^{n}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}\right)^{2}\right)^{-s/2}=2^{d}\sum_{x\in X}\left(1+\lvert x\rvert^{2}\right)^{-s/2}$
is finite for all $s>d$. ∎
3.2.2. Position spectral triple generates K-homology
Towards the end of this section, we shall prove that the position spectral triple $\xi_{\Lambda}^{\mathrm{Roe}}$ induces an isomorphism
$\operatorname{KK}(\mathrm{C}\ell_{d,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda))\xrightarrow{\sim}\operatorname{KK}(\mathrm{C}\ell_{d,d},\mathbb{R})\simeq\mathbb{Z}.$
The proof is based on these observations.
  1. (1)
    If $\Lambda=\mathbb{Z}^{d}$, then the K-homology class of the position spectral triple $\xi_{\mathbb{Z}^{d}}^{\mathrm{Roe}}$ pulls back to Kasparov’s Dirac element for the “real” manifold $\mathbb{T}^{d}$ which induces an isomorphism $\operatorname{KK}(\mathrm{C}\ell_{d,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d}))\xrightarrow{\sim}\mathbb{Z}$;
  2. (2)
    Any Delone set $\Lambda\subseteq\mathbb{R}^{d}$ is coarsely equivalent to $\mathbb{Z}^{d}$, hence $\xi^{\mathrm{Roe}}_{\Lambda}$ induces the same map as $\xi^{\mathrm{Roe}}_{\mathbb{Z}^{d}}$ up to an isomorphism of the K-theory groups
    $\operatorname{KK}(\mathrm{C}\ell_{d,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d}))\simeq\operatorname{KK}(\mathrm{C}\ell_{d,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)).$
Lemma 3.26.
The position spectral triple for $\Lambda=\mathbb{Z}^{d}$
(3.27) $\xi^{\mathrm{Roe}}_{\mathbb{Z}^{d}}\mathrel{\vcentcolon=}\left(\mathcal{C}_{\mathrm{Roe}}(\mathbb{Z}^{d})\otimes\mathrm{C}\ell_{0,d},\quad\ell^{2}(\mathbb{Z}^{d},\mathcal{K})\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d},\quad\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}\right),$
induces an isomorphism
$\operatorname{KK}(\mathrm{C}\ell_{d,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d}))\xrightarrow{\sim}\operatorname{KK}(\mathrm{C}\ell_{d,d},\mathbb{R})\simeq\mathbb{Z}.$
The proof below comes from [Ewert-Meyer:Coarse_geometry]*Section 4, which we recall for the reader’s convenience. Here we must use “real” C*-algebras and KKR-theory to describe the Dirac element of the “real” manifold $\mathbb{T}^{d}$. While passing to the “real” group C*-algebras of $\mathbb{Z}^{d}$ by Fourier transform, these $\operatorname{KKR}$-theory groups simply to $\operatorname{KKO}$-theory.
Proof.
We claim that the following diagram commutes:
[…]
Here the horizontal arrows are induced by inclusion maps
$\mathrm{C}_{0}(\mathbb{R}^{0,d})\hookrightarrow\mathrm{C}(\mathbb{T}^{d}),\qquad\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}}\hookrightarrow\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d})_{\mathbb{R}}.$
The first map comes from an inclusion of “real” manifolds $\mathbb{R}^{0,d}\hookrightarrow\mathbb{T}^{d}$. The map $\alpha_{\mathbb{R}^{0,d}}$ is the generator of the Bott periodicity of the “real” manifold $\mathbb{R}^{0,d}$, mapping a generator of $\operatorname{KKR}(\mathbb{C},\mathrm{C}_{0}(\mathbb{R}^{0,d})\otimes\mathbb{C}\ell_{0,d})$ to a generator of $\operatorname{KO}_{0}(\mathbb{R})\simeq\mathbb{Z}$, which is unique up to a sign. The vertical isomorphism $\operatorname{KKR}(\mathbb{C},\mathrm{C}(\mathbb{T}^{d})\otimes\mathbb{C}\ell_{0,d})$ is the composition
$\operatorname{KKR}(\mathbb{C},\mathrm{C}(\mathbb{T}^{d})\otimes\mathbb{C}\ell_{0,d})\xrightarrow{\operatorname{Ad}_{U^{*}}}\operatorname{KKR}(\mathbb{C},\mathrm{C}^{*}(\mathbb{Z}^{d})\otimes\mathbb{C}\ell_{0,d})\\ \xrightarrow{\sim}\operatorname{KKO}(\mathbb{R},\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}}\otimes\mathrm{C}\ell_{0,d})\xrightarrow{\sim}\operatorname{KKO}(\mathrm{C}\ell_{d,0},\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}}).$
The map $\alpha_{\mathbb{T}^{d}}$ is given by Kasparov’s Dirac element of the “real” manifold $(\mathbb{T}^{d},\mathfrak{r}\colon z\mapsto\overline{z})$, represented by the the Hodge–de Rham operator $\mathrm{d}+\mathrm{d}^{*}$. It follows from Lemma 2.8 that, via the Fourier transform isomorphism (2.7), the Hodge–de Rham operator $\mathrm{d}+\mathrm{d}^{*}$ is mapped to the operator ${\scriptstyle\mathsf{X}}=\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}$. We may further replace $\operatorname{KKR}(\mathrm{C}^{*}(\mathbb{Z}^{d})\otimes\mathbb{C}\ell_{d,0},\mathbb{C})$ by $\operatorname{KKO}(\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}}\otimes\mathrm{C}\ell_{d,0},\mathbb{R})$. This maps $\alpha_{\mathbb{T}^{d}}$ to the class represented by the real spectral triple
(3.29) $\left(\mathcal{A}\otimes\mathrm{C}\ell_{0,d},\quad\ell^{2}(\mathbb{Z}^{d})_{\mathbb{R}}\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d},\quad\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}\right),$
where $\mathcal{A}\subseteq\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}}$ is a dense, *-subalgebra. It is the pullback of the class of $\xi^{\mathrm{Roe}}_{\mathbb{Z}^{d}}$ in $\operatorname{KK}(\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)\otimes\mathrm{C}\ell_{0,d},\mathbb{R})$ of the spectral triple (3.27). This shows that the right part of the diagram commutes.
The left triangle of the diagram commutes, as was proven in [Ewert-Meyer:Coarse_geometry]*Proposition 11, because the inclusion $\mathbb{R}^{0,d}\hookrightarrow\mathbb{T}^{d}$ preserves the involution $\tau\colon z\mapsto-z$ of $\mathbb{T}^{d}$, thus gives an ideal inclusion $\mathrm{C}_{0}(\mathbb{R}^{0,d},\mathbb{C}\ell(\mathbb{R}^{0,d}))\hookrightarrow\mathrm{C}(\mathbb{T}^{d},\mathbb{C}\ell(\mathbb{T}^{d}))$. This ideal inclusion maps the Dirac element of $\mathbb{R}^{0,d}$ to the Dirac element of $\mathbb{T}^{d}$ via the Kasparov product.
Finally, both $\operatorname{KKR}(\mathbb{C},\mathrm{C}_{0}(\mathbb{R}^{0,d})\otimes\mathbb{C}\ell_{0,d})$ and $\operatorname{KKO}(\mathrm{C}\ell_{d,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d})_{\mathbb{R}})$ are isomorphic to the abelian group $\mathbb{Z}$. Thus the commutativity of the diagram claims that the image of $\xi_{\mathbb{Z}^{d}}^{\mathrm{Roe}}$ must be the image of $\alpha_{\mathbb{T}^{d}}$. This gives a surjective group homomorphism $\operatorname{KK}(\mathrm{C}\ell_{d,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d})_{\mathbb{R}})\to\mathbb{Z}$. Therefore, it follows from the K-theory of $\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d})_{\mathbb{R}}$ in (3.21) that this map must be an isomorphism. ∎
Lemma 3.30.
Let $L_{1},L_{2}\subseteq\mathbb{R}^{d}$ be Delone sets and set $L_{0}\mathrel{\vcentcolon=}L_{1}\cup L_{2}$.
  • For $i\in\{0,1,2\}$, let $\mathcal{H}_{i}\mathrel{\vcentcolon=}\ell^{2}(L_{i},\mathcal{K})$ be the standard ample $L_{i}$-modules of the discrete space $L_{i}$; and $\xi_{i}^{\mathrm{Roe}}\mathrel{\vcentcolon=}\xi^{\mathrm{Roe}}_{L_{i}}$ for the position spectral triples over $\mathrm{C}^{*}_{\mathrm{Roe}}(L_{i})$ as in (3.25).
  • For $i\in\{1,2\}$, let $\iota_{i}\colon L_{i}\hookrightarrow L_{0}$ be the isometric embeddings, and $V_{i}\colon\mathcal{H}_{i}\to\mathcal{H}_{0}$ be the isometries between Hilbert spaces induced by $\iota_{i}$, that is, $V_{i}\lvert x\rangle\mathrel{\vcentcolon=}\lvert\iota_{i}(x)\rangle$ for all $x\in L_{i}$.
Then the followings hold:
  1. (1)
    $V_{i}^{\prime}s$ are covering isometries for $\iota_{i}$. Hence $\operatorname{Ad}_{V_{i}}\colon T\mapsto V_{i}TV_{i}^{*}$ maps $\mathrm{C}^{*}_{\mathrm{Roe}}(L_{i})$ into $\mathrm{C}^{*}_{\mathrm{Roe}}(L)$, and induce isomorphisms
    $\iota_{i*}\colon\operatorname{KK}(\mathrm{C}\ell_{k,0},\mathrm{C}^{*}_{\mathrm{Roe}}(L_{i}))\to\operatorname{KK}(\mathrm{C}\ell_{k,0},\mathrm{C}^{*}_{\mathrm{Roe}}(L)).$
  2. (2)
    The following diagram commutes:
[…]
Proof.
(1) follows from Proposition 3.19 as both $L_{1}$ and $L_{2}$ are coarsely dense in $\mathbb{R}^{d}$, hence in $L_{0}$. In particular, we have a direct sum decomposition of Hilbert spaces
$\mathcal{H}_{0}=\mathcal{H}_{1}\oplus\mathcal{H}_{2},$
where $\mathcal{H}_{i}$ is identified with the image of the isometry $V_{i}\colon\mathcal{H}_{i}\hookrightarrow\mathcal{H}_{0}$.
Now we prove (2). Let $i\in\{0,1,2\}$. The position spectral triples $\xi^{\mathrm{Roe}}_{i}$ for $i\in\{0,1,2\}$ are given by the position operators $\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}^{i}_{j}\otimes\gamma^{j}$, where ${\scriptstyle\mathsf{X}}_{j}^{i}$ is the $j$-th position operator on $\mathcal{H}_{i}$:
${\scriptstyle\mathsf{X}}^{i}_{j}f(x)\mathrel{\vcentcolon=}x_{j}f(x),\quad x\in L_{i};$
The operators ${\scriptstyle\mathsf{X}}^{0}_{j}$ is diagonal for the decomposition $\mathcal{H}_{0}=\mathcal{H}_{1}\oplus\mathcal{H}_{2}$. Its restriction to $\mathcal{H}_{i}$ is ${\scriptstyle\mathsf{X}}^{i}_{j}$. Therefore, we have
${\scriptstyle\mathsf{X}}^{0}_{j}=V_{1}{\scriptstyle\mathsf{X}}_{j}^{1}V_{1}^{*}\oplus V_{2}{\scriptstyle\mathsf{X}}_{j}^{2}V_{2}^{*}=\operatorname{Ad}_{V_{1}}({\scriptstyle\mathsf{X}}_{j}^{1})\oplus\operatorname{Ad}_{V_{2}}({\scriptstyle\mathsf{X}}_{j}^{2}),$
and hence
$\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}^{0}_{j}\otimes\gamma^{j}=\operatorname{Ad}_{V_{1}}\left(\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}^{1}\otimes\gamma^{j}\right)\oplus\operatorname{Ad}_{V_{2}}\left(\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}^{2}\otimes\gamma^{j}\right).$
Therefore, the *-homomorphism $\operatorname{Ad}_{V_{1}}$ (resp. $\operatorname{Ad}_{V_{2}}$) pulls back the K-homology class represented by the position spectral triple $\xi_{0}^{\mathrm{Roe}}$ to the class represented by $\xi_{1}^{\mathrm{Roe}}$ (resp. $\xi_{2}^{\mathrm{Roe}}$ ). So the left (resp. right) triangle commutes by Lemma 2.13. ∎
Theorem 3.32.
For any Delone set $\Lambda$, the position spectral triple $\xi^{\mathrm{Roe}}_{\Lambda}$ induces an isomorphism
$\operatorname{KK}(\mathrm{C}\ell_{d,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda))\xrightarrow{\sim}\operatorname{KK}(\mathrm{C}\ell_{d,d},\mathbb{R})\simeq\mathbb{Z}.$
Proof.
Set $L_{1}=\mathbb{Z}^{d}$ and $L_{2}=\Lambda$. Then the position spectral triple $\xi^{\mathrm{Roe}}_{1}=\xi^{\mathrm{Roe}}_{\mathbb{Z}^{d}}$ induces an isomorphism
$\operatorname{KK}(\mathrm{C}\ell_{d,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d}))\xrightarrow{\sim}\operatorname{KK}(\mathrm{C}\ell_{d,d},\mathbb{R})\simeq\mathbb{Z}$
by Lemma 3.26. The isometry $V_{1}$ induces an isomorphism between the corresponding Kasparov groups by Lemma 3.30. Thus the commutativity of the left triangle in (3.31) implies that the map
$\operatorname{KK}(\mathrm{C}\ell_{d,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d}\cup\Lambda))\to\mathbb{Z}$
given by the position spectral triple $\xi^{\mathrm{Roe}}_{\mathbb{Z}^{d}\cup\Lambda}$, must be an isomorphism. Thus the position spectral triple $\xi_{2}^{\mathrm{Roe}}=\xi_{\Lambda}^{\mathrm{Roe}}$ must be an isomorphism as well, as it is a composition of isomorphisms
$\operatorname{KK}(\mathrm{C}\ell_{d,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda))\underset{\sim}{\xrightarrow{\iota_{1*}}}\operatorname{KK}(\mathrm{C}\ell_{d,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d}\cup\Lambda))\underset{\sim}{\xrightarrow{\xi_{0}^{\mathrm{Roe}}}}\mathbb{Z}.\qed$
4. Robustness of topological phases
Now we use the results from previous sections to compare the groupoid model $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$ with the coarse-geometric model $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$ of topological phases on an aperiodic lattice $\Lambda$. Topological phases described by the K-theory of $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$ are called strong in [Ewert-Meyer:Coarse_geometry]. We follow this terminology. Such phases are stable under perturbations which perserve the conjugate-linear and/or chiral symmtries of the system, are locally compact and controlled, and do not close the gap of the Hamiltonian. This is because such perturbations lift to homotopies in the Roe C*-algebra, and hence preserve K-theory.
We will explain which phases, described by the K-theory of $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$, have the same robustness. This is done by mapping $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$ into a Roe C*-algebra $\mathrm{C}^{*}_{\mathrm{Roe}}(\omega)$ using the localised regular representations $\pi_{\omega}$ depending on a choice of $\omega\in\Omega$, whereas all of these $\omega$ yield isomorphic Roe C*-algebras as they are Delone sets in $\mathbb{R}^{d}$. We also explain that “stacked” topological phases, coming from lower-dimensional Delone sets, are always weak. Both results can be viewed as generalisations of [Ewert-Meyer:Coarse_geometry]*Section 4, in which the Delone set $\Lambda$ is the periodic square lattice $\mathbb{Z}^{d}$.
4.1. Position spectral triples detect strong topological phases
The results of Theorem 3.9 and Theorem 3.32 allows us to compare the groupoid model and the coarse-geometric model on the level of both C*-algebras and K-theory (index pairing). The main theorem is the following:
Theorem 4.1 (Position spectral triples detect strong topological phases).
For every $\omega\in\Omega_{0}$, The following diagram commutes:
[…]
where:
  • $\pi_{\omega}^{N}\colon\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\otimes\mathbb{M}_{N}(\mathbb{R})\to\mathbb{B}(\ell^{2}(\omega,\mathcal{K}))$ is the entrywise extension of the localised regular representation at $\omega\in\Omega_{0}$ as in (3.7);
  • $\prescript{}{d}{\lambda}_{\Omega_{0}}$ is the bulk cycle as in (3.8);
  • $\xi^{\mathrm{Gpd}}_{\omega,N}$ is the spectral triple over $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\otimes\mathbb{M}_{N}(\mathbb{R})$ defined in (3.10); $\xi^{\mathrm{Roe}}_{\omega}$ is the spectral triple over $\mathrm{C}^{*}_{\mathrm{Roe}}(\omega)$ defined in (3.25);
  • all arrows are given by taking (unbounded) Kasparov products.
Proof.
We claim that $\pi^{N}_{\omega}\colon\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\otimes\mathbb{M}_{N}(\mathbb{R})\to\mathbb{B}(\ell^{2}(\omega,\mathcal{K}))$ maps into $\mathrm{C}^{*}_{\mathrm{Roe}}(\omega)$ defined by the standard ample $\omega$-module $\ell^{2}(\omega,\mathcal{K})$. Whenever this holds, then it follows that the K-homology class represented by the spectral triple (3.10)
$\xi_{\omega,N}^{\mathrm{Gpd}}\mathrel{\vcentcolon=}\left(\mathbb{M}_{N}(\mathrm{C}_{\mathrm{c}}(\mathcal{G}_{\Lambda}))\otimes\mathrm{C}\ell_{0,d},\quad\ell^{2}(\omega,\mathcal{K})\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d},\quad\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}\right)$
equals the pullback along $\pi^{N}_{\omega}$ of the K-homology class represented by (3.25):
$\xi^{\mathrm{Roe}}_{\omega}\mathrel{\vcentcolon=}\left(\mathcal{C}_{\mathrm{Roe}}(\omega)\otimes\mathrm{C}\ell_{0,d},\quad\ell^{2}(\omega,\mathcal{K})\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d},\quad\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}\right).$
We have shown in Theorem 3.9 that $\xi^{\mathrm{Gpd}}_{\omega,N}$ is a locally compact position spectral triple. This implies that $\pi_{\omega}(\mathrm{C}_{\mathrm{c}}(\mathcal{G}_{\Lambda})\otimes\mathbb{M}_{N}(\mathbb{R}))\subseteq\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$ by Theorem 3.23. Since $\mathrm{C}_{\mathrm{c}}(\mathcal{G}_{\Lambda})\otimes\mathbb{M}_{N}(\mathbb{R})$ is dense in $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\otimes\mathbb{M}_{N}(\mathbb{R})$, we conclude that the image of $\pi_{\omega}^{N}$ is contained in $\mathrm{C}^{*}_{\mathrm{Roe}}(\omega)$. Then it follows from the functoriality of the Kasparov product in Lemma 2.13 that the bottom left triangle commutes.
The commutativity of the top right triangle follows from the construction of $\xi^{\mathrm{Gpd}}_{\omega}$ in Theorem 3.9. Therefore, we conclude that the entire diagram commutes. ∎
Example 4.3.
Consider the periodic square lattice $\Lambda=\mathbb{Z}^{d}$. Then $\Omega_{\Lambda}\simeq\mathbb{T}^{d}$ and $\Omega_{0}=\{\omega\}$ is a singleton, in which $\omega$ can be chosen to be any point in $\Omega_{\Lambda}$ (cf. [Bourne-Mesland:Aperiodic]*Example 2.7). There is a unique evaluation point $\operatorname{ev}_{\omega}\colon\{\omega\}\to\mathbb{C}$, which gives an isomorphism $\operatorname{K}_{0}(\mathbb{C})\xrightarrow{\sim}\mathbb{Z}$. The spectral triple $\xi^{\mathrm{Gpd}}_{\omega,N}$ now becomes a spectral triple over $\mathrm{C}^{*}(\{\omega\}\rtimes\mathbb{Z}^{d})\otimes\mathbb{M}_{N}(\mathbb{R})\simeq\mathbb{M}_{N}(\mathrm{C}^{*}(\mathbb{Z}^{d}))$, given by the exterior product of (3.29) with the imprimivitity bimodule $(\mathbb{M}_{N}(\mathbb{R}),\mathcal{K},0)$ induced by any rank-$N$ corner embedding $\mathbb{R}^{N}\hookrightarrow\mathcal{K}$. Thus we have recovered the commutative diagram (3.28) in [Ewert-Meyer:Coarse_geometry]*Theorem 7.
Example 4.4 (Kubota’s construction).
If $\Lambda$ is aperiodic, i.e. there is no non-zero $x\in\mathbb{R}^{d}$ such that $\Lambda=\Lambda-x$, then Kubota has constructed in [Kubota:Controlled_topological_phases]*Lemma 2.24, a surjective groupoid homomorphism $\mathcal{G}^{\mathrm{coa}}_{\Lambda}\to\mathcal{G}_{\Lambda}$, from the coarse groupoid $\mathcal{G}^{\mathrm{coa}}_{\Lambda}$ (cf. [Skandalis-Tu-Yu:Coarse_groupoids]) to the groupoid of Delone sets, which induces an inclusion $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\to\mathrm{C}^{*}_{\mathrm{u,Roe}}(\Lambda)$ into the uniform Roe C*-algebra of $\Lambda$, and hence into $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$ because $\mathrm{C}^{*}_{\mathrm{u,Roe}}(\Lambda)\subseteq\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$. We claim that his construction indeed works for all $\omega\in\Omega_{0}$, and coincides with the representation $\pi_{\omega}$ as in (3.6).
Let $\Lambda$ be an aperiodic Delone set. Then there exists an open, dense embedding $\varphi_{\Lambda}\colon\Lambda\hookrightarrow\Omega_{0}$ sending $x$ to $\Lambda-x$. Therefore, $\Omega_{0}$ is a compactification of $\Lambda$. It follows from the universal property of the Stone–Čech compactification that there exists a unique surjection $\overline{\varphi}_{\Lambda}\colon\beta\Lambda\to\Omega_{0}$ extending $\varphi_{\Lambda}$.
The coarse groupoid $\mathcal{G}^{\mathrm{coa}}_{\Lambda}$ is the union of all $\overline{\mathcal{E}}$, where $\mathcal{E}$ is a controlled subset of $\Lambda\times\Lambda$ (cf. Definition 3.13), and the closure is taken inside the Stone–Čech compactification of $\Lambda\times\Lambda$. Let $\mathcal{E}\subseteq\Lambda\times\Lambda$ be controlled, then the function
$\mathcal{E}\to\mathbb{R}^{d},\quad(x,y)\mapsto y-x$
is bounded, hence extends to $\beta\mathcal{E}$. Therefore, there is a surjective groupoid homomorphism $\mathcal{G}^{\mathrm{coa}}_{\Lambda}\to\mathcal{G}_{\Lambda}$, which uniquely extends the map
$\Lambda\times\Lambda\to\mathcal{G}_{\Lambda},\quad(x,y)\mapsto(\varphi_{\Lambda}(x),y-x).$
This induces an injective *-homomorphism
$\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\to\mathrm{C}^{*}_{\mathrm{u,Roe}}(\Lambda).$
Kubota’s construction holds for all $\omega\in\Omega_{0}$, disregarding whether or not it is aperiodic. (Even if $\Lambda$ is aperiodic, an element $\omega\in\Omega_{0}$ can still be periodic for some $x\in\mathbb{R}^{d}$.) To see this, note that
$\varphi_{\omega}\colon\omega\to\Omega_{0},\qquad x\mapsto\omega-x$
is a continuous map into a compact Hausdorff space. It fails to be injective if $\Lambda$ is not aperiodic, but still extends uniquely to a continuous map $\overline{\varphi}_{\omega}\colon\beta\omega\to\Omega_{0}$. Thus Kubota’s construction gives a (not necessarily surjective) groupoid homomorphism $\mathcal{G}_{\omega}^{\mathrm{coa}}\to\mathcal{G}_{\Lambda}$, and hence a *-homomorphism $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\to\mathrm{C}^{*}(\mathcal{G}^{\mathrm{coa}}_{\omega})\simeq\mathrm{C}^{*}_{\mathrm{u,Roe}}(\omega)$. Composing with a rank-one corner embedding yields a map into $\mathrm{C}^{*}_{\mathrm{u,Roe}}(\omega)\otimes\mathbb{K}\subseteq\mathrm{C}^{*}_{\mathrm{Roe}}(\omega)$.
On the other hand, the *-homomorphism $\pi^{1}_{\omega}\colon\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\to\mathbb{B}(\ell^{2}(\omega,\mathcal{K}))$ in (3.6) maps $f\in\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$ to the operator $(\pi_{\Lambda}(f)_{x,y})_{x,y\in\Lambda}$, whose matrix elements are given by
$\pi_{\omega}(f)_{x,y}=\left\langle x\,\middle|\pi_{\omega}(f)\middle|\,\mathopen{}y\right\rangle=f(\omega-x,y-x)=f(\varphi_{\omega}(x),y-x).$
Therefore, it is induced by Kubota’s groupoid homomorphism $\mathcal{G}^{\mathrm{coa}}_{\omega}\to\mathcal{G}_{\Lambda}$.
4.2. Stacked topological phases are weak
It was explained in [Ewert-Meyer:Coarse_geometry] why certain topological phases described by $\operatorname{K}_{*}(\mathrm{C}^{*}(\mathbb{Z}^{d}))$ are “weak”. Let $\varphi\colon\mathbb{Z}^{d}\to\mathbb{Z}^{d+1}$ is an injective group homomorphism. It induces a map $\varphi_{*}\colon\mathrm{C}^{*}(\mathbb{Z}^{d})\to\mathrm{C}^{*}(\mathbb{Z}^{d+1})$ and hence maps
$\varphi_{*}\colon\operatorname{K}_{*}(\mathrm{C}^{*}(\mathbb{Z}^{d}))\to\operatorname{K}_{*}(\mathrm{C}^{*}(\mathbb{Z}^{d+1}))$
in K-theory. Topological phases that belong to the image of $\varphi_{*}$ can be thought of as “stacked” from the lower-dimensional lattice $\mathbb{Z}^{d}$. It was shown in [Ewert-Meyer:Coarse_geometry]*Proposition 10 that such phases are killed by the map $\operatorname{K}_{*}(\mathrm{C}^{*}(\mathbb{Z}^{d+1}))\to\operatorname{K}_{*}(\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d+1}))$ induced by the inclusion $\mathrm{C}^{*}(\mathbb{Z}^{d+1})\hookrightarrow\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d+1})$. The proof is based on the fact that $\varphi$ also induces a map between the Roe C*-algebras of $\mathbb{Z}^{d}$ and $\mathbb{Z}^{d+1}$, which factors through a flasque space if $\varphi$ is injective. Then this map induces zero maps in K-theory.
We shall explain in this section how this observation can be generalised to aperiodic lattices and higher-dimensional cases. That is, we consider Delone sets of the form $\Lambda\times L$, where $L$ is another Delone set. If $L$ has dimension $1$, then we may think of $\Lambda\times L$ as “stacking” $\Lambda$ along the direction of $L$. This also gives a *-homomorphism groupoid *-homomorphism $\varphi^{\mathrm{Gpd}}\colon\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\to\mathrm{C}^{*}(\mathcal{G}_{\Lambda\times L})$, which induces maps
$\varphi^{\mathrm{Gpd}}_{*}\colon\operatorname{KK}(\mathrm{C}\ell_{k,0},\mathrm{C}^{*}(\mathcal{G}_{\Lambda}))\to\operatorname{KK}(\mathrm{C}\ell_{k,0},\mathrm{C}^{*}(\mathcal{G}_{\Lambda\times L}))$
for all $k$. Such maps can be interpreted as “stacking” topological phases living on $\Lambda$ along the direction $L$. We will show that such topological phases are always weak, in the sense that they vanish in the K-theory of Roe C*-algebras.
We fix some notations. Let $\Lambda\subseteq\mathbb{R}^{m}$ be an $(r_{\Lambda},R_{\Lambda})$-Delone set and $L\subseteq\mathbb{R}^{n}$ be an $(r_{L},R_{L})$-Delone set. We write $\Omega_{\Lambda\times L}$, $\Omega_{\Lambda}$ and $\Omega_{L}$ for the closure of the orbit of $\Lambda\times L$, $\Lambda$ and $L$. To distinguish, we write
[…]
for corresponding abstract transversals.
Proposition 4.5.
The set
$\Lambda\times L\mathrel{\vcentcolon=}\left\{(x,a)\;\middle|\;x\in\Lambda,a\in L\right\}\subseteq\mathbb{R}^{m+n}$
is also a Delone set.
Proof.
Choose $0<r<R$ satisfying $r<\min\{r_{L},r_{\Lambda}\}$ and $R>\sqrt{R_{\Lambda}^{2}+R_{L}^{2}}$. Then for all $(x,a)\in\mathbb{R}^{m}\times\mathbb{R}^{n}$, we have
$\operatorname{B}((x,a),r)\subseteq\operatorname{B}(x,r_{\Lambda})\times\operatorname{B}(a,r_{L}),\quad\operatorname{B}((x,a),R)\supseteq\operatorname{B}(x,R_{\Lambda})\times\operatorname{B}(a,R_{L}).$
Hence
#$\#(\operatorname{B}((x,a),r)\cap(\Lambda\times L))$$\leq\#((\operatorname{B}(x,r_{\Lambda})\times\operatorname{B}(a,r_{L}))\cap(\Lambda\times L))$
=#$=\#((\operatorname{B}(x,r_{\Lambda})\cap\Lambda)\times(\operatorname{B}(a,r_{L})\cap L))$
=#$=\#(\operatorname{B}(x,r_{\Lambda})\cap\Lambda)\cdot\#(\operatorname{B}(a,r_{L})\cap L)$
$\leq 1,$
#$\#(\operatorname{B}((x,a),R)\cap(\Lambda\times L))$$\geq\#((\operatorname{B}(x,R_{\Lambda})\times\operatorname{B}(a,R_{L}))\cap(\Lambda\times L))$
=#$=\#((\operatorname{B}(x,R_{\Lambda})\cap\Lambda)\times(\operatorname{B}(a,R_{L})\cap L))$
=#$=\#(\operatorname{B}(x,R_{\Lambda})\cap\Lambda)\cdot\#(\operatorname{B}(a,R_{L})\cap L)$
$\geq 1.$
So $\Lambda\times L$ is an $(r,R)$-Delone set. ∎
A convenient way to define the “stacking” map is by describing the C*-algebra of the product Delone set as a minimal tensor product. We note the following well-known lemma:
Lemma 4.6.
Let $\mathcal{G}_{1}\rightrightarrows\Omega_{1}$ and $\mathcal{G}_{2}\rightrightarrows\Omega_{2}$ be étale groupoids. Let $\mathcal{G}_{1}\times\mathcal{G}_{2}$ be the product groupoid of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$. Then there is an isomorphism
$\Phi\colon\mathrm{C}^{*}(\mathcal{G}_{1})\otimes\mathrm{C}^{*}(\mathcal{G}_{2})\xrightarrow{\sim}\mathrm{C}^{*}(\mathcal{G}_{1}\times\mathcal{G}_{2}),\quad\Phi(f_{1}\otimes f_{2})(\gamma_{1},\gamma_{2})\mathrel{\vcentcolon=}f_{1}(\gamma_{1})f_{2}(\gamma_{2}).$
where $\otimes$ refers to the minimal tensor product.
Corollary 4.7.
There is a canonical isomorphism
$\mathrm{C}^{*}(\mathcal{G}_{\Lambda\times L})\simeq\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\otimes\mathrm{C}^{*}(\mathcal{G}_{L}).$
Proof.
By Lemma 4.6, it suffices to prove that the étale groupoids $\mathcal{G}_{\Lambda\times L}$ and $\mathcal{G}_{\Lambda}\times\mathcal{G}_{L}$ are isomorphic.
We claim that $\Omega_{\Lambda\times L}\simeq\Omega_{\Lambda}\times\Omega_{L}$ are isomorphic as $\mathbb{R}^{m+n}$-spaces. We have
$\Lambda\times L\in\operatorname{Del}_{(r_{\Lambda},R_{\Lambda})}(\mathbb{R}^{m})\times\operatorname{Del}_{(r_{L},R_{L})}(\mathbb{R}^{n}).$
The translation action of $\mathbb{R}^{m+n}=\mathbb{R}^{m}\times\mathbb{R}^{n}$ preserves $\operatorname{Del}_{(r_{\Lambda},R_{\Lambda})}(\mathbb{R}^{m})\times\operatorname{Del}_{(r_{L},R_{L})}(\mathbb{R}^{n})$. By Proposition 3.1, both $\operatorname{Del}_{(r_{\Lambda},R_{\Lambda})}(\mathbb{R}^{m})\times\operatorname{Del}_{(r_{L},R_{L})}(\mathbb{R}^{n})$ are closed, hence their product is a closed subset of $\mathcal{M}(\mathbb{R}^{m+n})$. Since the product of the orbits of $\Lambda$ and $L$ is contained in $\operatorname{Del}_{(r_{\Lambda},R_{\Lambda})}(\mathbb{R}^{m})\times\operatorname{Del}_{(r_{L},R_{L})}(\mathbb{R}^{n})$, we conclude that $\Omega_{\Lambda\times L}$, $\Omega_{\Lambda}\times\Omega_{L}$ are contained in it. Hence, $\Omega_{\Lambda\times L}\subseteq\Omega_{\Lambda}\times\Omega_{L}$ as it is the smallest closed subset in $\operatorname{Del}_{(r_{\Lambda},R_{\Lambda})}(\mathbb{R}^{m})\times\operatorname{Del}_{(r_{L},R_{L})}(\mathbb{R}^{n})$, which contains the product of the joint orbits of $\Lambda$ and $L$.
We claim that $\Omega_{\Lambda\times L}\supseteq\Omega_{\Lambda}\times\Omega_{L}$ as well. Let $(\omega,\ell)\in\Omega_{\Lambda}\times\Omega_{L}$, then there exists nets $(x_{\alpha})_{\alpha\in A}\subseteq\mathbb{R}^{m}$ and $(a_{\beta})_{\beta\in B}\subseteq\mathbb{R}^{n}$ such that
[…]
Then the net $(x_{\alpha},a_{\beta})_{\alpha\times\beta\in A\times B}$, where $A\times B$ carries the lexicographic order, satisfies
[…]
Then we conclude that $\Omega_{\Lambda\times L}\simeq\Omega_{\Lambda}\times\Omega_{L}$ as $\mathbb{R}^{m+n}$-spaces.
As a consequence, the action groupoids $\Omega_{\Lambda\times L}\rtimes\mathbb{R}^{m+n}$ and $(\Omega_{\Lambda}\times\Omega_{L})\rtimes\mathbb{R}^{m+n}$ are isomorphic topological groupoids, and have homeomorphic abstract transversals $\Omega_{0}^{\Lambda\times L}$ and $\Omega_{0}^{\Lambda}\times\Omega_{0}^{L}$. This implies an isomorphism of étale groupoids
$\mathcal{G}_{\Lambda\times L}\simeq\mathcal{G}_{\Lambda}\times\mathcal{G}_{L},\quad(\mu,z)\mapsto((\mu_{\Lambda},x),(\mu_{L},a))$
where $\mu_{\Lambda}$ and $\mu_{L}$ are the images of $\mu$ under the coordinate projections $\mathbb{R}^{m}\times\mathbb{R}^{n}\to\mathbb{R}^{m}$ and $\mathbb{R}^{m}\times\mathbb{R}^{n}\to\mathbb{R}^{n}$. This finishes the proof. ∎
Corollary 4.7 allows us to define a *-homomorphism between groupoid C*-algebras. Since $\mathcal{G}_{L}$ is an étale groupoid, its C*-algebra $\mathrm{C}^{*}(\mathcal{G}_{L})$ has a unit, given by the constant function $1_{\Omega_{0}^{L}}$ on the unit space. Define
(4.8) $\varphi^{\mathrm{Gpd}}\colon\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\to\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\otimes\mathrm{C}^{*}(\mathcal{G}_{L}),\quad\varphi^{\mathrm{Gpd}}(f)\mathrel{\vcentcolon=}f\otimes 1_{\Omega_{0}^{L}}.$
Then $\varphi^{\mathrm{Gpd}}$ gives a *-homomorphism $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\to\mathrm{C}^{*}(\mathcal{G}_{\Lambda\times L})$ under the isomorphism $\mathrm{C}^{*}(\mathcal{G}_{\Lambda\times L})\simeq\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\otimes\mathrm{C}^{*}(\mathcal{G}_{L})$ in Corollary 4.7. We write $\varphi^{\mathrm{Gpd}}_{N}\colon\mathbb{M}_{N}(\mathrm{C}^{*}(\mathcal{G}_{\Lambda}))\to\mathbb{M}_{N}(\mathrm{C}^{*}(\mathcal{G}_{\Lambda\times L}))$ for its entrywise extension to matrix algebras.
Now we pass to Roe C*-algebras. Let $\omega\in\Omega_{0}^{\Lambda}$ and $\ell\in\Omega_{0}^{L}$. Define their Roe C*-algebras $\mathrm{C}^{*}_{\mathrm{Roe}}(\omega)$ and $\mathrm{C}^{*}_{\mathrm{Roe}}(\omega\times\ell)$ using their standard ample modules $\ell^{2}(\omega,\mathcal{K})$ and $\ell^{2}(\omega\times\ell,\mathcal{K})$. Let
(4.9) $\varphi^{\mathrm{Roe}}\colon\mathbb{B}(\ell^{2}(\omega,\mathcal{K}))\to\mathbb{B}(\ell^{2}(\omega\times\ell,\mathcal{K})),\quad T\mapsto T\otimes\operatorname{id}_{\ell^{2}(\ell)}.$
Then
$\left\langle x,a\,\middle|\varphi^{\mathrm{Roe}}(T)\middle|\,\mathopen{}y,b\right\rangle=\left\langle x\,\middle|T\middle|\,\mathopen{}y\right\rangle\cdot\left\langle a\mid b\right\rangle=T_{x,y}\cdot\delta_{a,b}.$
So $\varphi^{\mathrm{Roe}}(T)$ is locally compact or controlled iff $T$ is locally compact or controlled (cf. Example 3.15). Thus $\varphi^{\mathrm{Roe}}$ maps $\mathrm{C}^{*}_{\mathrm{Roe}}(\omega)$ into $\mathrm{C}^{*}_{\mathrm{Roe}}(\omega\times\ell)$.
Lemma 4.10.
The map $\varphi^{\mathrm{Roe}}\colon\mathrm{C}^{*}_{\mathrm{Roe}}(\omega)\to\mathrm{C}^{*}_{\mathrm{Roe}}(\omega\times\ell)$ induces zero maps
$\varphi^{\mathrm{Roe}}_{*}\colon\operatorname{KK}(\mathrm{C}\ell_{k,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\omega))\to\operatorname{KK}(\mathrm{C}\ell_{k,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\omega\times\ell))$
for all $k$.
Proof.
Let
+:$\ell_{+}\mathrel{\vcentcolon=}\ell\cap(\mathbb{R}^{n-1}\times\mathbb{R}_{\geq 0}),\quad\ell_{-}\mathrel{\vcentcolon=}\ell\cap(\mathbb{R}^{n-1}\times\mathbb{R}_{<0}),$
then there are isometries
$V_{\pm}\colon\ell^{2}(\omega\times\ell_{\pm},\mathcal{K})\to\ell^{2}(\omega\times\ell,\mathcal{K}),$
which induce a diagonal embedding
$\mathbb{B}(\ell^{2}(\omega\times\ell_{+},\mathcal{K}))\oplus\mathbb{B}(\ell^{2}(\omega\times\ell_{-},\mathcal{K}))$ $\to\mathbb{B}(\ell^{2}(\omega\times\ell,\mathcal{K})),$
$(T_{+},T_{-})$ $\mapsto\operatorname{Ad}_{V_{+}}T_{+}+\operatorname{Ad}_{V_{-}}T_{-}.$
Let $T\in\mathbb{B}(\ell^{2}(\omega,\mathcal{K}))$. Then $T\otimes\operatorname{id}_{\ell^{2}(\ell)}=\operatorname{Ad}_{V_{+}}(T\otimes\operatorname{id}_{\ell^{2}(\ell_{+})})+\operatorname{Ad}_{V_{-}}(T\otimes\operatorname{id}_{\ell^{2}(\ell_{-})})$. Therefore, the map $\varphi^{\mathrm{Roe}}$ agrees with the following composition:
$\mathbb{B}(\ell^{2}(\omega,\mathcal{K}))\to\mathbb{B}(\ell^{2}(\omega\times\ell_{+},\mathcal{K}))\oplus\mathbb{B}(\ell^{2}(\omega\times\ell_{-},\mathcal{K}))\to\mathbb{B}(\ell^{2}(\omega\times\ell,\mathcal{K})),\\ T\mapsto\left(T\otimes\operatorname{id}_{\ell^{2}(\ell_{+})},T\otimes\operatorname{id}_{\ell^{2}(\ell_{-})}\right)\mapsto\operatorname{Ad}_{V_{+}}\left(T\otimes\operatorname{id}_{\ell^{2}(\ell_{+})}\right)+\operatorname{Ad}_{V_{-}}\left(T\otimes\operatorname{id}_{\ell^{2}(\ell_{-})}\right).$
A similar argument as the lines below (4.9) shows that $T\mapsto(T\otimes\operatorname{id}_{\ell^{2}(\ell_{+})},T\otimes\operatorname{id}_{\ell^{2}(\ell_{-})})$ maps $\mathrm{C}^{*}_{\mathrm{Roe}}(\omega)$ into $\mathrm{C}^{*}_{\mathrm{Roe}}(\omega\times\ell_{+})\oplus\mathrm{C}^{*}_{\mathrm{Roe}}(\omega\times\ell_{-})$. Both spaces $\omega\times\ell_{\pm}$ are flasque (cf. [Roe:Index_theory_coarse_geometry]*Definition 9.3). Hence, $\mathrm{C}^{*}_{\mathrm{Roe}}(\omega\times\ell_{\pm})$ have vanishing K-theory by an Eilenberg swindle argument, cf. [Roe:Index_theory_coarse_geometry]*Theorem 9.4. Thus $\varphi^{\mathrm{Roe}}_{*}$ vanishes as it factors through zero. ∎
Theorem 4.11 (Stacked topological phases are weak).
Let $\xi^{\mathrm{Gpd}}_{\omega\times\ell,N}$ be the spectral triple over $\mathrm{C}^{*}(\mathcal{G}_{\Lambda\times L})\otimes\mathbb{M}_{N}(\mathbb{R})$ defined in (3.10). Then its induced map
$\operatorname{KK}(\mathrm{C}\ell_{m+n,0},\mathrm{C}^{*}(\mathcal{G}_{\Lambda\times L}))\to\mathbb{Z}$
vanishes on the image of
$\varphi^{\mathrm{Gpd}}_{*}\colon\operatorname{KK}(\mathrm{C}\ell_{m+n},\mathrm{C}^{*}(\mathcal{G}_{\Lambda}))\to\operatorname{KK}(\mathrm{C}\ell_{m+n},\mathrm{C}^{*}(\mathcal{G}_{\Lambda\times L})).$
Proof.
We claim that the following diagram commutes:
${\mathbb{M}_{N}\mathrm{C}^{*}(\mathcal{G}_{\Lambda})}$${\mathbb{M}_{N}\mathrm{C}^{*}(\mathcal{G}_{\Lambda\times L})}$${\mathrm{C}^{*}_{\mathrm{Roe}}(\omega)}$${\mathrm{C}^{*}_{\mathrm{Roe}}(\omega\times\ell).}$$\scriptstyle{\varphi^{\mathrm{Gpd}}_{N}}$$\scriptstyle{\pi^{N}_{\omega}}$$\scriptstyle{\pi^{N}_{\omega\times\ell}}$$\scriptstyle{\varphi^{\mathrm{Roe}}}$
To see this, let $f\in\mathrm{C}_{\mathrm{c}}(\mathcal{G}_{\Lambda})$ and $S\in\mathbb{M}_{N}(\mathbb{R})$. It follows from (4.8) and (4.9) that:
$\left\langle x,a\,\middle|\varphi^{\mathrm{Roe}}\pi_{\omega}^{N}(f\otimes S)\middle|\,\mathopen{}y,b\right\rangle$ $=\left\langle x\,\middle|\pi_{\omega}(f)\middle|\,\mathopen{}y\right\rangle\cdot\delta_{a,b}\cdot S$
$=f(\omega-x,y-x)\cdot\delta_{a,b}\cdot S;$
$\left\langle x,a\,\middle|\pi_{\omega\times\ell}^{N}\varphi^{\mathrm{Gpd}}(f\otimes S)\middle|\,\mathopen{}y,b\right\rangle$ $=\varphi^{\mathrm{Gpd}}(f)(\omega\times\ell-(x,a),(y-x,b-a))\cdot S$
$=f(\omega-x,y-x)\cdot 1_{\Omega_{0}^{L}}(\ell-a,b-a)\cdot S$
$=f(\omega-x,y-x)\cdot\delta_{a,b}\cdot S$
hold for all $x,y\in\omega$ and $a,b\in\ell$. Therefore the maps $\varphi^{\mathrm{Roe}}\circ\pi_{\omega}^{N}$ coincides with $\pi_{\omega\times\ell}^{N}\circ\varphi^{\mathrm{Gpd}}$ on all elements of the form $f\otimes S$, and hence for $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\otimes\mathbb{M}_{N}(\mathbb{R})$. That is, the diagram above commutes.
The commutative diagram above gives a commutative diagram in K-theory. This, together with the the bottom left triangle of (4.2), gives the following commutative diagram:
${\operatorname{KK}(\mathrm{C}\ell_{m+n,0},\mathrm{C}^{*}(\mathcal{G}_{\Lambda}))}$${\operatorname{KK}(\mathrm{C}\ell_{m+n,0},\mathrm{C}^{*}(\mathcal{G}_{\Lambda\times L}))}$${\mathbb{Z}}$${\operatorname{KK}(\mathrm{C}\ell_{m+n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\omega))}$${\operatorname{KK}(\mathrm{C}\ell_{m+n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\omega\times\ell))}$$\scriptstyle{\varphi^{\mathrm{Gpd}}_{*}}$$\scriptstyle{\pi_{\omega}}$$\scriptstyle{\pi_{\omega\times\ell}}$$\scriptstyle{\xi^{\mathrm{Gpd}}_{\omega\times\ell}}$$\scriptstyle{\varphi^{\mathrm{Roe}}_{*}}$$\scriptstyle{\xi^{\mathrm{Roe}}_{\omega\times\ell}}$
where $\xi^{\mathrm{Gpd}}_{\omega\times\ell}$ and $\xi^{\mathrm{Roe}}_{\omega\times\ell}$ are the corresponding position spectral triples over $\mathrm{C}^{*}(\mathcal{G}_{\Lambda\times L})$ and $\mathrm{C}^{*}_{\mathrm{Roe}}(\omega\times\ell)$. It follows that the composition map $\xi^{\mathrm{Gpd}}_{\omega\times\ell}\circ\varphi_{*}=\xi^{\mathrm{Roe}}_{\omega\times\ell}\circ\varphi^{\mathrm{Roe}}_{*}\circ\pi_{\omega}$ factors through
$\varphi^{\mathrm{Roe}}_{*}\colon\operatorname{KK}(\mathrm{C}\ell_{m+n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\omega))\to\operatorname{KK}(\mathrm{C}\ell_{m+n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\omega\times\ell)),$
which vanishes by Lemma 4.10. Therefore $\xi^{\mathrm{Gpd}}_{\omega\times\ell}\circ\varphi^{\mathrm{Gpd}}_{*}$ must be the zero map. ∎
Example 4.12.
Let $\Lambda=\mathbb{Z}^{d}$ and $L=\mathbb{Z}$. Then $\mathcal{G}_{\Lambda}\simeq\mathbb{Z}^{d}$ and $\mathcal{G}_{L}\simeq\mathbb{Z}$ as topological groupoids. The unit spaces $\Omega_{0}^{\Lambda}$ and $\Omega_{0}^{\Lambda\times L}$ are both singletons. So there are unique localised regular representations $\pi\colon\mathrm{C}^{*}(\mathbb{Z}^{d})\to\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d})$ and $\mathrm{C}^{*}(\mathbb{Z}^{d+1})\to\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d+1})$. Let $N=1$. The *-homomorphism $\varphi^{\mathrm{Gpd}}_{1}\colon\mathrm{C}^{*}(\mathbb{Z}^{d})\to\mathrm{C}^{*}(\mathbb{Z}^{d+1})$ maps an element in $\mathrm{C}^{*}(\mathbb{Z}^{d+1})$ to its restriction to the first $d$-coordinates. If we identify $\mathrm{C}^{*}(\mathbb{Z}^{d})$ with $\mathrm{C}(\mathbb{T}^{d})$ via Fourier transform, then the map $\varphi^{\mathrm{Gpd}}$ sends $f\in\mathrm{C}(\mathbb{T}^{d})$ to $f\otimes 1\in\mathrm{C}(\mathbb{T}^{d})\otimes\mathrm{C}(\mathbb{T})=\mathrm{C}(\mathbb{T}^{d+1})$. The map $\varphi^{\mathrm{Gpd}}$ is induced by the injective group homomorphism
$\mathbb{Z}^{d}\to\mathbb{Z}^{d+1},\quad x\mapsto(x,0),$
hence its image belongs to the kernel of the map
$\operatorname{KK}(\mathrm{C}\ell_{k,0},\mathrm{C}^{*}(\mathbb{Z}^{d}))\to\operatorname{KK}(\mathrm{C}\ell_{k,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d}))$
by [Ewert-Meyer:Coarse_geometry]*Proposition 10.